### v1

# Lambda Expected Shortfall

Abstract
The Lambda Value-at-Risk ($\Lambda$-$\mathrm{VaR}$) is a generalization of the Value-at-Risk (VaR), which has been actively studied in quantitative finance. Over the past two decades, the Expected Shortfall (ES) has become one of the most important risk measures alongside VaR because of its various desirable properties in the practice of optimization, risk management, and financial regulation. Analogously to the intimate relation between ES and VaR, we introduce the Lambda Expected Shortfall ($\Lambda$-$\mathrm{ES}$), as a generalization of ES and a counterpart to $\Lambda$-$\mathrm{VaR}$. Our definition of $\Lambda$-$\mathrm{ES}$ has an explicit formula and many convenient properties, and we show that it is the smallest quasi-convex and law-invariant risk measure dominating $\Lambda$-$\mathrm{VaR}$ under mild assumptions. We examine further properties of $\Lambda$-ES, its dual representation, and related optimization problems.
Keywords: Lambda Value-at-Risk, quantiles, Expected Shortfall, quasi-convexity, dual representation
1 Introduction
In the landscape of quantitative finance and actuarial science, efficient and robust measurement of risk is paramount. Financial institutions and regulators make use of sophisticated tools to quantify potential losses and manage financial exposures effectively. Among the most widely adopted risk measures are the Value-at-Risk (VaR) and the Expected Shortfall (ES), each with distinct theoretical properties and practical implications. VaR has long served as a standard for risk assessment due to its intuitive interpretability. However, its well-documented limitations, such as the lack of subadditivity and non-convexity for general loss distributions and inability of capturing tail risk (see e.g., Daníelsson et al., 2001; McNeil et al., 2015; Embrechts et al., 2018), have spurred the development of more robust alternatives. ES, also known as the Conditional Value-at-Risk (CVaR), emerged as the most popular alternative, favored for its coherence (Artzner et al., 1999; Acerbi and Tasche, 2002), convexity (Föllmer and Schied, 2002; Frittelli and Rosazza Gianin, 2002), optimization properties (Rockafellar and Uryasev, 2002; Embrechts et al., 2022), and axiomatization via portfolio concentration (Wang and Zitikis, 2021), although it suffers from the lack of elicitability (Gneiting, 2011; Ziegel, 2016; Kou and Peng, 2016; Fissler and Ziegel, 2016).
As a flexible generalization of VaR, the class of Lambda Value-at-Risk ($\Lambda$-VaR) was introduced by Frittelli et al. (2014). The class of $\Lambda$-VaR offers enhanced adaptability for modeling diverse risk preferences and regulatory contexts beyond a fixed confidence level. $\Lambda$-VaR is found to satisfy several useful properties in finance, including monotonicity, cash-subadditivity, elicitability (Bellini and Bignozzi, 2015), robustness (Burzoni et al., 2017), and quasi-star-shapeness (Han et al., 2025). Bellini and Peri (2022) obtained an axiomatic characterization of $\Lambda$-VaR, in particular justifying the choice of $\Lambda$ to be a (weakly) decreasing function. As a risk measure, $\Lambda$-VaR has also been studied from practical aspects such as estimation and backtesting (Hitaj et al., 2018; Corbetta and Peri, 2018), distributionally robust optimizations (Han and Liu, 2025), capital allocations (Ince et al., 2022; Liu, 2025), and optimal insurance problems (Boonen et al., 2025). While $\Lambda$-VaR successfully broadens the scope of VaR, it retains the essential drawbacks of VaR for not being convex and not being able to capture tail risk. A natural remedy for the problem is to introduce an equally flexible generalization of ES as an alternative to $\Lambda$-VaR. However, a suitable way of defining such a risk measure that preserves its desirable properties and strong theoretical foundations has not been found.
This paper addresses this gap by introducing the Lambda Expected Shortfall ($\Lambda$-ES), a novel and theoretically sound risk measure designed to serve as the natural counterpart to $\Lambda$-VaR. One may argue that there are many ways to generalize ES to a class of risk measures parametrized by a function $\Lambda$. A key consideration in ES and its generalization is its consistency with respect to portfolio diversification, modelled via convexity by Föllmer and Schied (2002); Frittelli and Rosazza Gianin (2002). For general risk measures, Cerreia-Vioglio et al. (2011) argued that diversification preferences should be modelled by quasi-convexity, which is equivalent to convexity for monetary risk measures. Keeping this property as our fundamental requirement for a generalization of ES, we find that there is one formulation that has the most advantages, inspired by a recent $\Lambda$-VaR representation result of Han et al. (2025), explained below. For a decreasing function $\Lambda$, we define $\Lambda$-ES of a random variable $X$ by
$\sup_{x\in\mathbb{R}}\left(\mathrm{ES}_{\Lambda(x)}(X)\wedge x\right).$ (1)
We demonstrate that, among other potential candidate definitions, $\Lambda$-ES defined as (1) possesses several critical properties (Proposition 2), analogous to those that establish ES as an improved alternative to VaR. Specifically, we prove that, under a mild assumption on $\Lambda$, $\Lambda$-ES defined in (1) is the smallest quasi-convex and law-invariant risk measure that dominates $\Lambda$-VaR (Theorem 2). This generalization is a significant theoretical advancement, extending the dominance relationship from VaR versus ES (Delbaen, 2012; Föllmer and Schied, 2016) to the more flexible $\Lambda$-VaR versus $\Lambda$-ES framework. As a by-product, we obtain a new result on the domination of ES over VaR (Theorem 1) that is stronger than several classic results in the literature.
Beyond the foundational definition and properties of $\Lambda$-ES, which are the topics of Section 3, we proceed to conduct a comprehensive analysis of this new class of risk measures. In Section 4, we obtain a dual representation of $\Lambda$-ES (Theorem 3), offering deeper insights into its theoretical structure and connections to quasi-convex cash-subadditive risk measures. In Section 5, we explore the properties of $\Lambda$-ES in optimization problems, both as an objective function to minimize and as a constraint to impose, and analyze various forms of convexity in relevant reformulations of ES optimization problems. As standard in the risk measures literature, the main results are formulated on the space $L^{\infty}$ of essentially bounded random variables. In Section 6, results are naturally extended to the space $L^{1}$ of integrable random variables, sometimes under slightly stronger assumptions. Section 7 concludes the paper. Some alternative potential formulations for $\Lambda$-ES are discussed in Appendix A, demonstrating why our proposed definition is the most robust, theoretically consistent, and desirable for risk management applications.
2 VaR, Lambda VaR and ES
2.1 Risk measures
Let $L^{0}$ be the space of all random variables on an atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$, $L^{1}$ be the space of all random variables with finite mean, and $L^{\infty}$ be the set of all essentially bounded random variables. Write $\overline{\mathbb{R}}=[-\infty,\infty]$ and $\mathbb{R}_{+}=[0,\infty)$. For any $n\in\mathbb{N}$, denote by $n]=\{1,\dots,n\}$. For any $x,y\in\overline{\mathbb{R}}$, write $x\wedge y=\min\{x,y\}$, $\vee y=\max\{x,y\}, x+=x$$_{+}=x\vee 0$, and $x_{-}=x\wedge 0$. For any function $f:\mathbb{R}\to\mathbb{R}$ and $x\in\mathbb{R}$, we write $f(x-)=\lim_{y\uparrow x}f(y)$ and $f(x+)=\lim_{y\downarrow x}f(y)$, if they exist.
We start with risk measures that are used to quantify risks. A risk measure is a mapping $\rho:\mathcal{X}\to\overline{\mathbb{R}}$, where $\mathcal{X}$ is the space of random variables where $\rho$ is defined. Below, we list several properties that a risk measure may satisfy. The functional $\rho$ is called a monetary risk measure if it satisfies
  • Monotonicity: $\rho(X)\geqslant\rho(Y)$ for all $X,Y\in\mathcal{X}$ and $X\geqslant Y$ almost surely;
  • Cash additivity (or translation invariance): $\rho(X+m)=\rho(X)+m$ for all $X\in\mathcal{X}$ and $m\in\mathbb{R}$.
As standard considerations for risk measures (e.g., Artzner et al., 1999), monotonicity naturally means that increasing the amount of loss will lead to more risk of the financial position, whereas cash additivity ensures that $\rho(X)$ is the total capital required to be added to the financial position to make it acceptable given the loss of $X\in\mathcal{X}$. It is also natural to assume a monetary risk measure satisfies
  • Normalization: $\rho(t)=t$ for all $t\in\mathbb{R}$.
A monetary risk measure is called coherent (Artzner et al., 1999) if it further satisfies [^1] Whenever convexity or subadditivity is discussed, the range of $\rho$ includes at most one of $\infty$ and -$\infty$ to avoid $\infty$-$\infty$.
  • Positive homogeneity: $\rho(\gamma X)=\gamma\rho(X)$ for all $X\in\mathcal{X}$ and $\gamma\in(0,\infty)$;
  • Subadditivity: $\rho(X+Y)\leqslant\rho(X)+\rho(Y)$ for all $X,Y\in\mathcal{X}$;
whereas a monetary risk measure is called a convex risk measure (Föllmer and Schied, 2002; Frittelli and Rosazza Gianin, 2002) if it further satisfies
  • Convexity: $\rho(\gamma X+(1-\gamma)Y)\leqslant\gamma\rho(X)+(1-\gamma)\rho(Y)$ for all $X,Y\in\mathcal{X}$ and $\gamma\in[0,1]$.
Convexity is motivated by diversification effects in risk measurement. We refer to Föllmer and Schied (2016) for a comprehensive review of properties of monetary risk measures. Beyond the monetary framework, El Karoui and Ravanelli (2009) argued that cash additivity fails to incorporate the ambiguity of interest rates and thus proposed the property of
  • Cash subadditivity: $\rho(X+m)\leqslant\rho(X)+m$ for all $X\in\mathcal{X}$ and $m\in\mathbb{R}_{+}$.
For cash-subadditive risk measures, Cerreia-Vioglio et al. (2011) argued that the diversification effect is characterized by
  • Quasi-convexity: $\rho(\gamma X+(1-\gamma)Y)\leqslant\max\{\rho(X),\rho(Y)\}$ for all $X,Y\in\mathcal{X}$ and $\gamma\in[0,1]$.
Many commonly used convex risk measures (such as the Expected Shortfall defined below) also satisfy law invariance and concavity with respect to distribution mixtures, defined below. Let $\mathcal{M}_{c}(\mathbb{R})$ denote the set of compactly supported distributions on $\mathbb{R}$.
  • Law-invariance: $\rho(X)=\rho(Y)$ for all $X,Y\in\mathcal{X}$ with the same distribution.
  • Concavity (resp. quasi-concavity) in mixtures: $\rho$ is law-invariant and the function $F\mapsto\rho(X_{F})$ on $\mathcal{M}_{c}(\mathbb{R})$ is concave (resp. quasi-concave), where $X_{F}$ is a random variable with distribution $F\in\mathcal{M}_{c}(\mathbb{R})$.
Further properties of risk measures that we will consider in this paper include
  • SSD-consistency: $\rho(X)\geqslant\rho(Y)$ for all $X,Y\in\mathcal{X}$ and $X\succeq_{\rm icx}Y$. [^2] Here, SSD represents second-order stochastic dominance. For $X,Y\in\mathcal{X}$, we say that $X$ dominates $Y$ in increasing convex order, denoted by $X\succeq_{\rm icx}Y$, if $\mathbb{E}[f(X)]\geqslant\mathbb{E}[f(Y)]$ for all increasing and convex functions $f:\mathbb{R}\to\mathbb{R}$.
  • $L^{1}$-continuity: $\rho(X_{n})\to\rho(X)$ for all $X,X_{1},X_{2}\dots\in\mathcal{X}$ and $X_{n}\stackrel{{\scriptstyle L^{1}}}{{\to}}X$ as $n\to\infty$.
For the financial interpretation and characterization of the properties introduced above, we refer to Föllmer and Schied (2002, convexity), Cerreia-Vioglio et al. (2011, quasi-convexity), Mao and Wang (2020, SSD-consistency), and Wang et al. (2020, concavity in mixtures).
2.2 VaR and Lambda VaR
The Value-at-Risk (VaR) at level $\alpha\in[0,1]$ is defined as the left-quantile, namely,
$\mathrm{VaR}_{\alpha}(X)=\inf\{x\in\mathbb{R}:\mathbb{P}(X\leqslant x)\geqslant\alpha\},~~~X\in L^{0}.$
Note that for all $X\in L^{0}$, we have $\mathrm{VaR}_{0}(X)=$-$\infty$ and $\mathrm{VaR}_{1}(X)$ is the essential supremum of $X$. Similarly, the upper Value-at-Risk ($\mathrm{VaR}^{+}$) at level $\alpha\in[0,1]$ is defined as the right-quantile:
$\mathrm{VaR}^{+}_{\alpha}(X)=\inf\{x\in\mathbb{R}:\mathbb{P}(X\leqslant x)>\alpha\},~~~X\in L^{0}.$
Using these formulations, $\mathrm{VaR}^{+}_{0}(X)$ is the essential infimum of $X$ and $\mathrm{VaR}^{+}_{1}(X)=\infty$. Moreover, $\mathrm{VaR}_{\alpha}(X),\mathrm{VaR}^{+}_{\alpha}(X)\in\mathbb{R}$ for any $\alpha\in(0,1)$. Both versions of VaR satisfy monotonicity, cash additivity, positive homogeneity, law invariance, and quasi-concavity in mixtures, but not quasi-convexity, concavity in mixtures, or SSD-consistency. Indeed, the latter three properties are equivalent for the class of distortion risk measures; see Wang et al. (2020, Theorem 3).
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function. Throughout, all terms like “increasing” and “decreasing” are in the weak sense. The $\Lambda$-Value-at-Risk ($\Lambda$-VaR, or $\Lambda$-quantile), denoted by $\mathrm{VaR}_{\Lambda}:L^{0}\to\overline{\mathbb{R}}$, is defined as
$\mathrm{VaR}_{\Lambda}(X)=\inf\{x\in\mathbb{R}:\mathbb{P}(X\leqslant x)\geqslant\Lambda(x)\}=\sup\{x\in\mathbb{R}:\mathbb{P}(X\leqslant x)<\Lambda(x)\},~~~X\in L^{0}.$
Frittelli et al. (2014) originally introduced $\Lambda$-VaR focusing on the case that $\Lambda$ is increasing. Bellini and Bignozzi (2015) showed that $\Lambda$-VaR with a decreasing $\Lambda$ satisfies elicitability, and it is not true for increasing $\Lambda$ (Burzoni et al., 2017). A more decisive result is the axiomatic justification of Bellini and Peri (2022) for using a decreasing function $\Lambda$. Han et al. (2025) further showed that $\Lambda$-VaR with a decreasing $\Lambda$ is cash-subadditive and hence $L^{\infty}$-continuous, but with an increasing $\Lambda$ even $L^{\infty}$-continuity fails. For these reasons, we focus on the case of decreasing $\Lambda$.
The quantile has two versions, and so does $\Lambda$-VaR. We define the upper $\Lambda$-Value-at-Risk ($\Lambda$-$\mathrm{VaR}^{+}$ or upper $\Lambda$-quantile), denoted by $\mathrm{VaR}^{+}_{\Lambda}:L^{0}\to\overline{\mathbb{R}}$, as
$\mathrm{VaR}^{+}_{\Lambda}(X)=\inf\{x\in\mathbb{R}:\mathbb{P}(X\leqslant x)>\Lambda(x)\}=\sup\{x\in\mathbb{R}:\mathbb{P}(X\leqslant x)\leqslant\Lambda(x)\},~~~X\in L^{0}.$
If $\Lambda$ is a constant $\alpha\in[0,1]$ then $\mathrm{VaR}_{\Lambda}=\mathrm{VaR}_{\alpha}$ (resp. $\mathrm{VaR}^{+}_{\Lambda}=\mathrm{VaR}^{+}_{\alpha}$), which is the left- (resp. right-) quantile at level $\alpha$. The $\Lambda$-VaR is monotone, but not cash additive or positively homogeneous, thus losing some usual properties of VaR.
2.3 Expected Shortfall
We present the standard risk measure in banking regulation, the Expected Shortfall (ES), via several different formulations. First, as the most standard definition, ES at level $\alpha\in[0,1]$ is formally defined as the mapping $\mathrm{ES}_{\alpha}:L^{\infty}\to\mathbb{R}$ given by
$\mathrm{ES}_{\alpha}(X)=\frac{1}{1-\alpha}\int_{\alpha}^{1}\mathrm{VaR}_{\beta}(X)\mathrm{d}\beta\mbox{ for $\alpha\in[0,1)$},$ (2)
and $\mathrm{ES}_{1}(X)=\mathrm{VaR}_{1}(X)$. Note that $\mathrm{ES}_{0}=\mathbb{E}$ and the definition of $\mathrm{ES}_{\alpha}$ in (2) can be easily extended to $L^{1}$, and we will discuss risk measures on $L^{1}$ in Section 6. An ES satisfies all properties listed in Section 2.1.
Second, as shown by Rockafellar and Uryasev (2002), for $\alpha\in[0,1]$, $\mathrm{ES}_{\alpha}$ can be equivalently formulated by
$\mathrm{ES}_{\alpha}(X)=\min_{x\in\mathbb{R}}\left\{x+\frac{1}{1-\alpha}\mathbb{E}[(X-x)_{+}]\right\},~~~X\in L^{\infty},$ (3)
where we use the convention $0/0=0$ and $x/0=\infty$ for $x>0$. The representation (3) connects to VaR via
$\operatorname*{arg\,min}_{x\in\mathbb{R}}\left\{x+\frac{1}{1-\alpha}\mathbb{E}[(X-x)_{+}]\right\}=\left\{\begin{array}[]{ll}[\mathrm{VaR}_{\alpha}(X),\mathrm{VaR}^{+}_{\alpha}(X)],&\mbox{if}~\alpha\in[0,1),\\ \mathrm{VaR}_{1}(X),&\mbox{if}~\alpha=1,\end{array}\right.~~~X\in L^{\infty}.$ (6)
We may call (3) and (6) the Rockafellar–Uryasev (RU) relation for VaR and ES.
Third, it is known that the risk measure $\mathrm{ES}_{\alpha}$ is the smallest law-invariant coherent risk measure dominating $\mathrm{VaR}_{\alpha}$ (Delbaen, 2012, Theorem 52). Convexity is important and relevant for risk management, and for this reason, ES is regarded as an improvement of VaR. In the next result, we show that $\mathrm{ES}_{\alpha}$ is also the smallest mapping dominating $\mathrm{VaR}_{\alpha}$ satisfying quasi-convexity and law invariance. As far as we know, this result is new, and it is based on a VaR-ES asymptotic equivalence result of Wang and Wang (2015) and a result in Embrechts et al. (2015) on the sum of negatively dependent sequences. Throughout the paper, we write $\rho\geqslant\widetilde{\rho}$ for mappings $\rho:\mathcal{X}\to\overline{\mathbb{R}}$ and $\widetilde{\rho}:\widetilde{\mathcal{X}}\to\overline{\mathbb{R}}$ to represent the dominance of $\rho$ over $\widetilde{\rho}$ on their common domain (i.e., $\rho(X)\geqslant\widetilde{\rho}(X)$ for all $X\in\mathcal{X}\cap\widetilde{\mathcal{X}}$), and typically we have either $X\subseteq\widetilde{\mathcal{X}}$ or $\widetilde{X}\subseteq\mathcal{X}$.
Theorem 1.
For any $\alpha\in(0,1]$,
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{\infty}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (7)
For any $\alpha\in[0,1)$,
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{\infty}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}^{+}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (8)
Proof.
(i) We first prove (7). Let $\rho$ be quasi-convex and law-invariant satisfying $\rho\geqslant\mathrm{VaR}_{\alpha}$. If $\alpha=1$, it is clear that (7) holds, because $\mathrm{VaR}_{1}$ is quasi-convex and law-invariant. Next, suppose $\alpha\in(0,1)$. For any $X\in L^{\infty}$ with distribution $F$, we have
$\rho(X)$ $=\sup\left\{\max\{\rho(X_{1}),\dots,\rho(X_{n})\}:{X_{i}\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}F,~i\in[n]}\right\}$ [law invariance of $\rho$]
$\geqslant\sup\left\{\rho\left(\frac{X_{1}+\dots+X_{n}}{n}\right):{X_{i}\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}F,~i\in[n]}\right\}$ [quasi-convexity of $\rho$]
$\geqslant\sup\left\{\mathrm{VaR}_{\alpha}\left(\frac{X_{1}+\dots+X_{n}}{n}\right):{X_{i}\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}F,~i\in[n]}\right\}$ $\rho\geqslant\mathrm{VaR}_{\alpha}$]
$=\frac{1}{n}\sup\left\{\mathrm{VaR}_{\alpha}\left({X_{1}+\dots+X_{n}}\right):{X_{i}\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}F,~i\in[n]}\right\}$ [positive homogeneity of $\mathrm{VaR}_{\alpha}$]
$\to\mathrm{ES}_{\alpha}(X),~~~\mbox{as $n\to\infty$}.$ [Corollary 3.7 of Wang and Wang (2015)]
This shows that $\rho\geqslant\mathrm{ES}_{\alpha}$ for any $\rho$ in the set in (7). Since $\mathrm{ES}_{\alpha}$ also satisfies law invariance and quasi-convexity, we know that the minimum of the set in (7) is $\mathrm{ES}_{\alpha}$.
(ii) For $\alpha\in(0,1)$, the result in part (i) implies that $\mathrm{ES}_{\alpha}$ is the smallest quasi-convex and law-invariant risk measure that dominates $\mathrm{VaR}_{\alpha}$, and since $\mathrm{ES}_{\alpha}\geqslant\mathrm{VaR}_{\alpha}^{+}\geqslant\mathrm{VaR}_{\alpha}$, the conclusion also holds for $\mathrm{VaR}_{\alpha}^{+}$.
For $\alpha=0$ and $X\in L^{\infty}$ with distribution $F$, write $M=\mathrm{VaR}_{1}(X)-\mathrm{VaR}^{+}_{0}(X)$. For any $n\in\mathbb{N}$, by Corollary A.3 of Embrechts et al. (2015), there exist $\widetilde{X}_{i}\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}F$, $i\in[n]$, such that
$\left|\frac{1}{n}\sum^{n}_{i=1}\widetilde{X}_{i}-\mathbb{E}[X]\right|\leqslant\frac{M}{n}.$
Hence,
$\frac{1}{n}\sum^{n}_{i=1}\widetilde{X}_{i}\geqslant\mathbb{E}[X]-\frac{M}{n}.$
It yields that
$\mathbb{E}[X]\geqslant\frac{1}{n}\mathrm{VaR}^{+}_{0}\left({\widetilde{X}_{1}+\dots+\widetilde{X}_{n}}\right)\geqslant\mathbb{E}[X]-\frac{M}{n}.$
Therefore,
$\frac{1}{n}\sup\left\{\mathrm{VaR}^{+}_{0}\left({X_{1}+\dots+X_{n}}\right):{X_{i}\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}F,~i\in[n]}\right\}\to\mathbb{E}[X],~~~\mbox{as }n\to\infty.$
Hence, we have $\rho(X)\geqslant\mathbb{E}[X]$ in the same sense as the argument in (i). As $\mathbb{E}$ dominates the essential infimum $\mathrm{VaR}^{+}_{0}$, it implies that $\mathbb{E}$ is the smallest quasi-convex and law-invariant risk measure that dominates $\mathrm{VaR}^{+}_{0}$. The proof is complete. ∎
Theorem 1 is stronger than two classic results: Föllmer and Schied (2016, Theorem 4.67), which requires $\rho$ in (7) to be convex, monetary and Fatou-continuous, and Delbaen (2012, Theorem 52), which requires $\rho$ in (7) to be coherent. [^3] We say a risk measure $\rho:\mathcal{X}\to\mathbb{R}$ is Fatou-continuous if it is lower semicontinuous under bounded pointwise convergence: For all bounded $X,X_{1},X_{2},\dots\in\mathcal{X}$ such that $X_{n}\to X$ pointwise as $n\to\infty$, $\rho(X)\leqslant\liminf_{n\to\infty}\rho(X_{n})$. Both of the two results above further assumed that $\rho$ takes finite values and $\alpha\in(0,1)$, but these differences are not essential.
We note that (7) fails for $\alpha=0$ because $\mathrm{VaR}_{0}=$-$\infty$ is quasi-convex, and the smallest quasi-convex and law-invariant risk measure dominating $\mathrm{VaR}_{0}$ is itself instead of $\mathrm{ES}_{0}=\mathbb{E}$. Similarly, (8) fails for $\alpha=1$ because $\mathrm{VaR}_{1}^{+}=\infty$ is quasi-convex, and the smallest quasi-convex and law-invariant risk measure dominating $\mathrm{VaR}_{1}^{+}$ is itself instead of $\mathrm{ES}_{1}=\mathrm{VaR}_{1}$.
For $\alpha\in(0,1)$, the statement in Theorem 1 extends directly to $L^{1}$, but for $\alpha=0$ and $\alpha=1$, some minor adjustments are needed, which we discuss in Section 6.
3 Lambda ES
The main question of this paper is how to define a counterpart to $\Lambda$-VaR that is similar to ES as a counterpart to VaR. We will call this counterpart $\Lambda$-Expected Shortfall, or $\Lambda$-ES for short.
3.1 Requirements for Lambda ES
By defining a new class of risk measures, there should be some clear gain. Otherwise, the newly defined class is not useful. The following properties are also satisfied by $\Lambda$-VaR, and they will be considered as basic requirements for $\Lambda$-ES. We believe their desirability is self-evident.
  • $\Lambda$-ES should be parameterized only by the function $\Lambda$.
  • $\Lambda$-ES should coincide with $\mathrm{ES}_{\alpha}$ when $\Lambda$ is equal to a constant $\alpha\in[0,1]$.
  • $\Lambda$-ES should increase as $\Lambda$ increases.
  • $\Lambda$-ES should be monotone and law invariant.
The next four properties are additional crucial requirements for $\Lambda$-ES to be considered a useful alternative to $\Lambda$-VaR, and they highlight the features of ES over VaR.
  • $\Lambda$-ES should dominate $\Lambda$-VaR. This is analogous to the dominance of ES over VaR.
  • $\Lambda$-ES should be quasi-convex. This should be the key improvement of $\Lambda$-ES over $\Lambda$-VaR so that it captures the diversification effects.
  • $\Lambda$-ES should be SSD-consistent. This property allows for $\Lambda$-ES to capture strong risk aversion in decision theory and to make consistent risk assessment.
  • $\Lambda$-ES should be $L^{1}$-continuous. This guarantees the robustness of $\Lambda$-ES with respect to slight changes to random losses, sharing the same advantage as ES and other coherent risk measures in risk management practice.
Some other properties, such as normalization, cash subadditivity, and quasi-concavity in mixtures are also desirable and natural from the corresponding properties of ES, but they may be less crucial as a criterion. They will all be satisfied by our proposed candidate for $\Lambda$-ES.
3.2 A formal definition of Lambda ES
For a decreasing function $\Lambda:\mathbb{R}\to[0,1]$, to define $\Lambda$-ES using the idea of Theorem 1, the formulation is obvious:
$\min\{\rho:L^{\infty}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}_{\Lambda}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (9)
Theorem 1 of Han et al. (2025) gives a representation result of $\Lambda$-VaR. Below, we restate the result and extend it to $\Lambda$-$\mathrm{VaR}^{+}$.
Proposition 1.
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function. The risk measures $\mathrm{VaR}_{\Lambda}$ and $\mathrm{VaR}^{+}_{\Lambda}$ admit the following representations.
$\mathrm{VaR}_{\Lambda}(X)=\sup_{x\in\mathbb{R}}\left(\mathrm{VaR}_{\Lambda(x)}(X)\wedge x\right)=\inf_{x\in\mathbb{R}}\left(\mathrm{VaR}_{\Lambda(x)}(X)\vee x\right),~~~X\in L^{0},$ (10)
$\mathrm{VaR}^{+}_{\Lambda}(X)=\sup_{x\in\mathbb{R}}\left(\mathrm{VaR}^{+}_{\Lambda(x)}(X)\wedge x\right)=\inf_{x\in\mathbb{R}}\left(\mathrm{VaR}^{+}_{\Lambda(x)}(X)\vee x\right),~~~X\in L^{0}.$ (11)
Proof.
Equation (10) holds directly by Theorem 1 of Han et al. (2025) for all decreasing functions $\Lambda:\mathbb{R}\to[0,1]$ that are not constantly 0. For $\Lambda\equiv 0$ (meaning that $\Lambda$ is constantly 0), we have
$\sup_{x\in\mathbb{R}}\left(-\infty\wedge x\right)=\inf_{x\in\mathbb{R}}\left(-\infty\vee x\right)=-\infty=\mathrm{VaR}_{0}(X).$
Thus (10) holds for all decreasing functions $\Lambda:\mathbb{R}\to[0,1]$.
The proof of (11) is in the same spirit as that in Theorem 1 of Han et al. (2025) to prove (10). We write it below for completeness. For a decreasing function $\Lambda:\mathbb{R}\to[0,1]$ and $X\in L^{0}$, we have
$\mathrm{VaR}^{+}_{\Lambda}(X)$ $=\sup\{x\in\mathbb{R}:\mathbb{P}(X\leqslant x)\leqslant\Lambda(x)\}$
$=\sup\{x\in\mathbb{R}:\mathrm{VaR}^{+}_{\Lambda(x)}(X)\geqslant x\}$
$=\sup\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\wedge x:\mathrm{VaR}_{\Lambda(x)}(X)\geqslant x\}\leqslant\sup_{x\in\mathbb{R}}\left\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\wedge x\right\},$
and similarly,
$\mathrm{VaR}^{+}_{\Lambda}(X)$ $=\sup\{x\in\mathbb{R}:\mathrm{VaR}^{+}_{\Lambda(x)}(X)\geqslant x\}$
$=\inf\{x\in\mathbb{R}:\mathrm{VaR}^{+}_{\Lambda(x)}(X)<x\}$
$=\inf\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\vee x:\mathrm{VaR}_{\Lambda(x)}(X)<x\}\geqslant\inf_{x\in\mathbb{R}}\left\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\vee x\right\}.$
Therefore, we have
$\mathrm{VaR}^{+}_{\Lambda}(X)\leqslant\sup_{x\in\mathbb{R}}\left\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\wedge x\right\}\leqslant\inf_{x\in\mathbb{R}}\left\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\vee x\right\}\leqslant\mathrm{VaR}^{+}_{\Lambda}(X).$
The proof is complete. ∎
Inspired by (10), we formally define the Lambda Expected Shortfall as follows.
Definition 1 ($\Lambda$-Expected Shortfall).
For a decreasing function $\Lambda:\mathbb{R}\to[0,1]$, the $\Lambda$-Expected Shortfall ($\Lambda$-ES) is defined as the risk measure $\mathrm{ES}_{\Lambda}:L^{\infty}\to\mathbb{R}$ given by
$\mathrm{ES}_{\Lambda}(X)=\sup_{x\in\mathbb{R}}\left(\mathrm{ES}_{\Lambda(x)}(X)\wedge x\right),~~~X\in L^{\infty}.$ (12)
Figure 1 illustrates the definition of $\Lambda$-ES when the function $\Lambda$ is continuous and discontinuous, respectively. Writing $x^{*}=\mathrm{ES}_{\Lambda}(X)$, we can see that $(x^{*},x^{*})$ is the unique intersection point between the graph (linearly interpolated) of the function $x\mapsto\mathrm{ES}_{\Lambda(x)}(X)$ and the graph of the identity. From the right panel of Figure 1, we can also see that whether $x\mapsto\mathrm{ES}_{\Lambda(x)}(X)$ is left- or right-continuous at $x^{*}$ (or neither) does not matter.
$x^{*}$$\mathrm{ES}_{\Lambda}(X)$$f_{1}(x)=\mathrm{ES}_{\Lambda(x)}(X)$$f_{2}(x)=x$$xx^{*}$$\mathrm{ES}_{\Lambda}(X)$$f_{1}(x)=\mathrm{ES}_{\Lambda(x)}(X)$$f_{2}(x)=x$$x$
Figure 1: Illustration of $\mathrm{ES}_{\Lambda}$ in Definition 1; left panel shows $\mathrm{ES}_{\Lambda}$ for a continuous $\Lambda$; right panel shows $\mathrm{ES}_{\Lambda}$ for $\Lambda$ that is discontinuous at $x^{*}=\mathrm{ES}_{\Lambda}(X)$
By definition, $\mathrm{ES}_{\Lambda}$ is finite on $L^{\infty}$; see also Proposition 2 (a) below. The next result shows that $\mathrm{ES}_{\Lambda}$ in Definition 1 satisfies (9), generalizing the fact that $\mathrm{ES}_{\alpha}$ satisfies (7). Moreover, it admits an alternative representation (13), which can be clearly seen from Figure 1.
Theorem 2.
The following statements hold.
  1. (i)
    For a decreasing function $\Lambda:\mathbb{R}\to(0,1]$, the smallest quasi-convex and law-invariant risk measure on $L^{\infty}$ dominating $\mathrm{VaR}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^∞→Rρ⩾VaR_Λ and $\rho$ is quasi-convex and law-invariant}.
  2. (ii)
    For a decreasing function $\Lambda:\mathbb{R}\to[0,1)$, the smallest quasi-convex and law-invariant risk measure on $L^{\infty}$ dominating $\mathrm{VaR}^{+}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^∞→Rρ⩾VaR^+_Λ and $\rho$ is quasi-convex and law-invariant}.
Moreover, the following identity holds for all decreasing functions $\Lambda:\mathbb{R}\to[0,1]$:
$\mathrm{ES}_{\Lambda}(X)=\inf_{x\in\mathbb{R}}\left(\mathrm{ES}_{\Lambda(x)}(X)\vee x\right),~~~X\in L^{\infty}.$ (13)
Proof.
Using (10)–(12), we can see that $\mathrm{ES}_{\Lambda}$ dominates $\mathrm{VaR}_{\Lambda}$ for $\Lambda:\mathbb{R}\to[0,1]$ (resp. $\mathrm{VaR}^{+}_{\Lambda}$ for $\Lambda:\mathbb{R}\to[0,1)$) because $\mathrm{ES}_{\alpha}\geqslant\mathrm{VaR}_{\alpha}$ for all $\alpha\in[0,1]$ (resp. $\mathrm{ES}_{\alpha}\geqslant\mathrm{VaR}^{+}_{\alpha}$ for all $\alpha\in[0,1)$). Moreover, $\mathrm{ES}_{\Lambda}$ is law-invariant by definition. Next, we show that $\mathrm{ES}_{\Lambda}$ is quasi-convex. Note that for any given $\alpha\in[0,1]$, $\mathrm{ES}_{\alpha}$ is quasi-convex. Further, an increasing transform of a quasi-convex function is quasi-convex, as well as the supremum of a set of quasi-convex functions. Using these facts,
$\mathrm{ES}_{\Lambda(x)}\mbox{ is quasi-convex for each $x\in\mathbb{R}$}$
$\implies\mathrm{ES}_{\Lambda(x)}\wedge x\mbox{ is quasi-convex for each $x\in\mathbb{R}$}$
$\implies\sup_{x\in\mathbb{R}}\left(\mathrm{ES}_{\Lambda(x)}\wedge x\right)\mbox{ is quasi-convex.}$
Therefore, $\mathrm{ES}_{\Lambda}$ is a quasi-convex and law-invariant risk measure dominating $\mathrm{VaR}_{\Lambda}$ for $\Lambda:\mathbb{R}\to[0,1]$ (resp. $\mathrm{VaR}^{+}_{\Lambda}$ for $\Lambda:\mathbb{R}\to[0,1)$).
Next, for $\Lambda:\mathbb{R}\to(0,1]$, we show that for any $\rho$ that is quasi-convex, law-invariant, and satisfying $\rho\geqslant\mathrm{VaR}_{\Lambda}$, it must be $\rho\geqslant\mathrm{ES}_{\Lambda}$. For any $X\in L^{\infty}$, we have:
$\rho(X)\geqslant\mathrm{VaR}_{\Lambda}(X)$ $\implies\rho(X)\geqslant\sup_{x\in\mathbb{R}}\left(\mathrm{VaR}_{\Lambda(x)}(X)\wedge x\right)$
$\implies\mbox{for all $x\in\mathbb{R}:~$}\rho(X)\geqslant\left(\mathrm{VaR}_{\Lambda(x)}(X)\wedge x\right)$
$\implies\mbox{for all $x\in\mathbb{R}:~$}\rho(X)\geqslant\mathrm{VaR}_{\Lambda(x)}(X)\mbox{ or }\rho(X)\geqslant x$
[Theorem 1] $\implies\mbox{for all $x\in\mathbb{R}:~$}\rho(X)\geqslant\mathrm{ES}_{\Lambda(x)}(X)\mbox{ or }\rho(X)\geqslant x$
$\implies\rho(X)\geqslant\sup_{x\in\mathbb{R}}\left(\mathrm{ES}_{\Lambda(x)}(X)\wedge x\right)=\mathrm{ES}_{\Lambda}(X).$
For $\Lambda:\mathbb{R}\to[0,1)$, for any $\rho$ that is quasi-convex, law-invariant, and satisfying $\rho\geqslant\mathrm{VaR}^{+}_{\Lambda}$, we have $\rho(X)\geqslant\mathrm{ES}_{\Lambda}(X)$ for any $X\in L^{\infty}$ with the same argument as above by replacing $\mathrm{VaR}_{\Lambda(x)}$ by $\mathrm{VaR}^{+}_{\Lambda(x)}$ for all $x\in\mathbb{R}$. This completes the proof of statements (i) and (ii). The final statement in (13) follows from (10), by noting that an ES curve $\alpha\mapsto\mathrm{ES}_{\alpha}(X)$ for $X\in L^{\infty}$ can be written as a VaR (resp. $\mathrm{VaR}^{+}$) curve $\alpha\mapsto\mathrm{VaR}_{\alpha}(Y)$ for some $Y\in L^{0}$ on $\alpha\in(0,1]$ (resp. $\alpha\in[0,1)$); see e.g., Lemma 4.5 of Burzoni et al. (2022). ∎
In Theorem 2, the reason to exclude 0 in part (i) and $1$ in part (ii) from the range of $\Lambda$ is the same as that in Theorem 1, where 0 (resp. $1$) is excluded from the domination of $\mathrm{ES}_{\alpha}$ over $\mathrm{VaR}_{\alpha}$ (resp. $\mathrm{VaR}_{\alpha}^{+}$).
An immediate consequence of (12)–(13) is that, for any $X\in L^{\infty}$ and $x\in\mathbb{R}$, we have
$\mathrm{ES}_{\Lambda(x+)}(X)\leqslant x\leqslant\mathrm{ES}_{\Lambda(x-)}(X)~\iff~\mathrm{ES}_{\Lambda}(X)=x.$ (14)
This is also illustrated in Figure 1. As a result of (14), for any $X\in L^{\infty}$ and $x\in\mathbb{R}$, we have
$\mathrm{ES}_{\Lambda(x)}(X)=x~\implies~\mathrm{ES}_{\Lambda}(X)=x;$ (15)
moreover, if $\Lambda$ is continuous, then
$\mathrm{ES}_{\Lambda(x)}(X)=x~\iff~\mathrm{ES}_{\Lambda}(X)=x.$ (16)
The relations (14)–(16) will be convenient in some proof arguments.
Remark 1.
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function and $\rho_{\Lambda}=\mathrm{VaR}_{\Lambda}$, $\mathrm{VaR}^{+}_{\Lambda}$ or $\mathrm{ES}_{\Lambda}$. The supremum in $\rho=\sup_{x\in\mathbb{R}}\{\rho_{\Lambda(x)}\wedge x\}$ is a maximum when $\Lambda$ is left-continuous; similarly the infimum in $\rho=\inf_{x\in\mathbb{R}}\{\rho_{\Lambda(x)}\vee x\}$ is a minimum when $\Lambda$ is right-continuous; see Figure 1 for an illustration.
It is clear that $\mathrm{ES}_{\Lambda}$ is parameterized only by the function $\Lambda$, and $\mathrm{ES}_{\Lambda}=\mathrm{ES}_{\alpha}$ when $\Lambda\equiv\alpha$ for some $\alpha\in(0,1)$. It satisfies several other desirable properties as a good candidate for $\Lambda$-$\mathrm{ES}$ as discussed in Section 3.1, which we summarize in the following result.
Proposition 2.
For any decreasing function $\Lambda:\mathbb{R}\to[0,1]$, the risk measure $\mathrm{ES}_{\Lambda}$ satisfies the following properties: (a) $\mathrm{ES}_{\Lambda}\geqslant\mathrm{ES}_{\Lambda^{\prime}}$ when $\Lambda\geqslant\Lambda^{\prime}$; (b) $\mathrm{ES}_{\Lambda}$ is monotone; (c) $\mathrm{ES}_{\Lambda}\geqslant\mathrm{VaR}_{\Lambda}$; (d) $\mathrm{ES}_{\Lambda}$ is quasi-convex; (e) $\mathrm{ES}_{\Lambda}$ is cash-subadditive; (f) $\mathrm{ES}_{\Lambda}$ is normalized; (g) $\mathrm{ES}_{\Lambda}$ is SSD-consistent; (h) $\mathrm{ES}_{\Lambda}$ is quasi-concave in mixtures; (i) $\mathrm{ES}_{\Lambda}$ is $L^{1}$-continuous when $\Lambda$ takes values in $[0,1)$.
Proof.
Items (a) and (b) are straightforward because $\mathrm{ES}_{\alpha}(X)$ is monotone (increasing) in both $\alpha$ and $X$, and the supremum of monotone transformations of $\mathrm{ES}_{\alpha}(X)$ is also monotone.
Items (c) and (d) are implied by Theorem 2.
To see item (e), for $c\in\mathbb{R}_{+}$ and $X\in L^{\infty}$, we have
$\mathrm{ES}_{\Lambda}(X+c)$ $=\sup_{x\in\mathbb{R}}\{\mathrm{ES}_{\Lambda(x)}(X+c)\wedge x\}$
$=\sup_{x\in\mathbb{R}}\{(\mathrm{ES}_{\Lambda(x)}(X)+c)\wedge x\}$ [cash additivity of ES]
$=\sup_{x\in\mathbb{R}}\{\mathrm{ES}_{\Lambda(x)}(X)\wedge(x-c)\}+c$
$\leqslant\sup_{x\in\mathbb{R}}\{\mathrm{ES}_{\Lambda(x)}(X)\wedge x\}+c=\mathrm{ES}_{\Lambda}(X)+c.$
Item (f) follows from (15).
Item (g) follows by applying Lemma 4 of Han et al. (2025), using the fact that $\mathrm{ES}_{\Lambda}$ is cash subadditive, monotone, quasi-convex, and law-invariant. Cash subadditivity is proved in item (e). Law invariance of $\mathrm{ES}_{\Lambda}$ is straightforward from the representation in (12) and the law invariance of ES.
For item (h), we first note that $\mathrm{ES}_{\alpha}$ is concave in mixtures (Wang et al., 2020, Theorem 3) for each $\alpha\in[0,1]$. Since quasi-concavity is preserved under increasing transforms, we know that $\mathrm{ES}_{\alpha}\vee x$ is also quasi-concave in mixtures. By using (13) and the fact that the infimum of quasi-concave functions is quasi-concave, we know that $\mathrm{ES}_{\Lambda}$ is quasi-concave in mixtures.
To prove item (i), first note that $\mathrm{ES}_{\alpha}$ is $L^{1}$-continuous (e.g., Rüschendorf, 2013, Corollary 7.10) for each $\alpha\in[0,1)$. Take any random variable $X$ and any sequence $(X_{n})_{n\in\mathbb{N}}$ in $L^{\infty}$ such that $X_{n}\to X$ in $L^{1}$ as $n\to\infty$. Let $f_{n}:x\mapsto\mathrm{ES}_{\Lambda(x)}(X_{n})-x$ and $f:x\mapsto\mathrm{ES}_{\Lambda(x)}(X)-x$. By (12), for any $y,z$ with $y<\mathrm{ES}_{\Lambda}(X)<z$, we have $f(y)>0>f(z)$. Therefore, because $f_{n}\to f$ pointwise, we have $f_{n}(y)>0>f_{n}(z)$ for $n$ large enough. This implies $y\leqslant\mathrm{ES}_{\Lambda}(X_{n})\leqslant z$ via (14). Since $y,z$ are arbitrarily close to $\mathrm{ES}_{\Lambda}(X)$, we know $\mathrm{ES}_{\Lambda}(X_{n})\to\mathrm{ES}_{\Lambda}(X)$. ∎
By item (a) of Proposition 2, it is straightforward that $\mathrm{ES}_{\Lambda}$ is bounded above by $\mathrm{ES}_{1}$ and below by $\mathbb{E}$ on $L^{\infty}$. This statement also holds when $\mathrm{ES}_{\Lambda}$ is formulated on larger spaces such as $L^{1}$.
The assumption that $\Lambda$ does not take the value $1$ in item (i) is not dispensable, noting that $\mathrm{ES}_{1}$ is not $L^{1}$-continuous.
We close this section with a result showing that although $\Lambda$-ES is quasi-convex, it is not convex and not concave in mixtures in general, unless it is an ES. This result also highlights the fact that quasi-convexity and convexity are different in strength for cash-subadditive risk measures, although they coincide for monetary risk measures, as shown by Cerreia-Vioglio et al. (2011).
Proposition 3.
For any decreasing function $\Lambda:\mathbb{R}\to[0,1]$, the following statements are equivalent.
  1. (i)
    The risk measure $\mathrm{ES}_{\Lambda}$ is convex.
  2. (ii)
    The risk measure $\mathrm{ES}_{\Lambda}$ is concave in mixtures.
  3. (iii)
    The function $\Lambda$ is constant on $\mathbb{R}$.
Proof.
“(iii) $\Rightarrow$ (i)” and “(iii) $\Rightarrow$ (ii)” follow from the facts that ES is convex and ES is concave in mixtures.
“(i) $\Rightarrow$ (iii)”: Suppose that $\mathrm{ES}_{\Lambda}$ is convex and for contradiction that $\Lambda$ is not constant on $\mathbb{R}$. There exist $x>y$ with $\Lambda(x-)<\Lambda((x+y)/2)\leqslant\Lambda(y)$. Take $X,Y\in L^{\infty}$ with $1-\mathbb{P}(X=x)=\mathbb{P}(X=y)=\Lambda((x+y)/2)$ and $Y=y$. It follows that $\mathrm{ES}_{\Lambda}(Y)=y$ and $\mathrm{ES}_{\Lambda}((X+Y)/2)=(x+y)/2$. Because $\Lambda(x-)<\Lambda((x+y)/2)$, we have $\mathrm{ES}_{\Lambda(x-)}(X)<x$. By (14), we have $\mathrm{ES}_{\Lambda}(X)<x$. It follows that $\mathrm{ES}_{\Lambda}(X)/2+\mathrm{ES}_{\Lambda}(Y)/2<\mathrm{ES}_{\Lambda}((X+Y)/2)$, contradicting the convexity of $\mathrm{ES}_{\Lambda}$. Therefore, $\Lambda$ is constant on $\mathbb{R}$.
“(ii) $\Rightarrow$ (iii)”: Suppose that $\Lambda$ is not constant on $\mathbb{R}$. Since $\Lambda$ is bounded, it cannot be concave. Hence, there exist distinct points $x,y,z\in\mathbb{R}$ and $\theta\in(0,1)$ such that $z=\theta x+(1-\theta)y$ and $\Lambda(z)<\theta\Lambda(x)+(1-\theta)\Lambda(y).$ By the continuity of linear functions, there exists $\gamma\in(0,1)$ in any neighborhood of $\theta$ such that $z<\gamma x+(1-\gamma)y$ and $\Lambda(z)<\gamma\Lambda(x)+(1-\gamma)\Lambda(y).$ Write $p=\Lambda(x)$, $q=\Lambda(y)$ and $r=\gamma p+(1-\gamma)q$.
Take independent events $A,B,C\in\mathcal{F}$ such that $\mathbb{P}(A)=1-p$, $\mathbb{P}(B)=1-q$, and $\mathbb{P}(C)=\gamma$. For some constant $K>\max\{-x,-y\}$ (to be determined later), let
$X=x\mathds{1}_{A}-K\mathds{1}_{A^{c}},~~Y=y\mathds{1}_{B}-K\mathds{1}_{B^{c}},\mbox{~~and~~}Z=\mathds{1}_{C}X+\mathds{1}_{C^{c}}Y.$
We can calculate $\mathrm{ES}_{p}(X)=x$ and $\mathrm{ES}_{q}(Y)=y$. By (15), we have $\mathrm{ES}_{\Lambda}(X)=x$ and $\mathrm{ES}_{\Lambda}(Y)=y$. Note that the distribution of $Z$ is the mixture of those of $X$ and $Y$ with weights $\gamma$ and $(1-\gamma)$ respectively. We will show $\mathrm{ES}_{\Lambda}(Z)\leqslant z$ for large $K$, which, together with $z<\gamma\mathrm{ES}_{\Lambda}(X)+(1-\gamma)\mathrm{ES}_{\Lambda}(Y)$, disproves the concavity in mixtures of $\mathrm{ES}_{\Lambda}$.
Note that $\mathbb{P}(Z=-K)=\gamma\mathbb{P}(X=-K)+(1-\gamma)\mathbb{P}(Y=-K)=\gamma p+(1-\gamma)q=r$ and $\Lambda(z)<r$. Therefore,
$\mathrm{ES}_{\Lambda(z)}(Z)=\frac{1}{1-\Lambda(z)}\left(-(r-\Lambda(z))K+{\int_{r}^{1}\mathrm{VaR}_{\beta}(Z)\mathrm{d}\beta}\right)\leqslant-K\frac{r-\Lambda(z)}{1-\Lambda(z)}+\frac{1-r}{1-\Lambda(z)}\max\{x,y\},$
which tends to -$\infty$ as $K\to\infty$. In particular, for some $K$ large enough, we have $\mathrm{ES}_{\Lambda(z)}(Z)\leqslant z$. Using (13), we get $\mathrm{ES}_{\Lambda}(Z)\leqslant z$. ∎
4 Dual representation
We now study the dual representation of $\mathrm{ES}_{\Lambda}$ as a quasi-convex and cash-subadditive risk measure, in the form of Cerreia-Vioglio et al. (2011). Denote by $\mathcal{M}_{1,f}=\mathcal{M}_{1,f}(\Omega,\mathcal{F},\mathbb{P})$ the set of all finitely additive probability measures that are absolutely continuous with respect to $\mathbb{P}$. The following result shows the dual representation of $\mathrm{ES}_{\Lambda}$ as a direct consequence of its definition in (12).
Theorem 3.
For any decreasing function $\Lambda:\mathbb{R}\to[0,1]$, the risk measure $\mathrm{ES}_{\Lambda}$ adopts the following representation:
$\mathrm{ES}_{\Lambda}(X)=\sup_{\mathbb{Q}\in\mathcal{M}_{1,f}}R(\mathbb{E}_{\mathbb{Q}}[X],\mathbb{Q}),~~~X\in L^{\infty},$ (17)
where for $(t,\mathbb{Q})\in\mathbb{R}\times\mathcal{M}_{1,f}$,
$R(t,\mathbb{Q})=\sup_{x\in\mathbb{R}}\left\{t\wedge x~:~\Lambda(x)\geqslant 1-\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}},~\mathbb{Q}\mbox{-almost surely}\right\},$ (18)
where we write $\mathrm{d}\mathbb{P}/\mathrm{d}\mathbb{Q}=1/(\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P})$ with $1/0=\infty$. Moreover, the following statements hold.
  1. (i)
    The supremum in (17) is a maximum if $\Lambda$ is left-continuous.
  2. (ii)
    $(t,\mathbb{Q})\mapsto R(t,\mathbb{Q})$ is upper semicontinuous, quasi-concave, and increasing in $t$;
  3. (iii)
    $\inf_{t\in\mathbb{R}}R(t,\mathbb{Q})=\inf_{t\in\mathbb{R}}R(t,\mathbb{Q}^{\prime})$ for all $\mathbb{Q},\mathbb{Q}^{\prime}\in\mathcal{M}_{1,f}$;
  4. (iv)
    $R(t_{1},\mathbb{Q})-R(t_{2},\mathbb{Q})\leqslant t_{1}-t_{2}$ for all $t_{1}\geqslant t_{2}$ and $\mathbb{Q}\in\mathcal{M}_{1,f}$.
Proof.
Define
$\mathcal{P}_{\Lambda(x)}=\left\{\mathbb{Q}\in\mathcal{M}_{1,f}~:~\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\leqslant\frac{1}{1-\Lambda(x)},~\mathbb{P}\mbox{-almost surely}\right\},~~x\in\mathbb{R}.$
For any $X\in L^{1}$ and $x\in\mathbb{R}$, we have
$\mathrm{ES}_{\Lambda}(X)$ $=\sup_{x\in\mathbb{R}}\left(\mathrm{ES}_{\Lambda(x)}(X)\wedge x\right)$ [Theorem 2]
$=\sup_{x\in\mathbb{R}}\left\{\max_{\mathbb{Q}\in\mathcal{P}_{\Lambda(x)}}\mathbb{E}_{\mathbb{Q}}[X]\wedge x\right\}$ [Theorem 4.52 of Föllmer and Schied (2016)]
$=\sup_{x\in\mathbb{R}}\max_{\mathbb{Q}\in\mathcal{P}_{\Lambda(x)}}\left\{\mathbb{E}_{\mathbb{Q}}[X]\wedge x\right\}$
$=\sup_{\mathbb{Q}\in\mathcal{M}_{1,f}}\sup_{x\in\mathbb{R}}\left\{\mathbb{E}_{\mathbb{Q}}[X]\wedge x~:~\Lambda(x)\geqslant 1-\frac{1}{\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}},~\mathbb{P}\mbox{-almost surely}\right\},$ [definition of $\mathcal{P}_{\Lambda(x)}$]
where the supremum over $x\in\mathbb{R}$ can be changed to a maximum when $\Lambda$ is left-continuous. This implies statement (i). Further because
$\mathbb{Q}\left(\Lambda(x)\geqslant 1-\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\right)=1$ $\iff\mathbb{E}_{\mathbb{P}}\left[\mathds{1}_{\left\{\Lambda(x)\geqslant 1-\frac{1}{\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}}\right\}}\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\right]=1$
$\iff\mathbb{P}\left(\left\{\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}=0~\mbox{or}~\Lambda(x)\geqslant 1-\frac{1}{\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}}\right\}\right)=1$
$\iff\mathbb{P}\left(\Lambda(x)\geqslant 1-\frac{1}{\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}}\right)=1,$
we have
$\mathrm{ES}_{\Lambda}(X)=\sup_{\mathbb{Q}\in\mathcal{M}_{1,f}}\sup_{x\in\mathbb{R}}\left\{\mathbb{E}_{\mathbb{Q}}[X]\wedge x~:~\Lambda(x)\geqslant 1-\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}},~\mathbb{Q}\mbox{-almost surely}\right\}.$
Now it remains to check statements (ii)-(iv).
(ii) Upper semicontinuity can be seen by showing that for all $t_{0}\in\mathbb{R}$,
$\limsup_{t\to t_{0}}R(t,\mathbb{Q})$ $=\lim_{\delta\downarrow 0}\sup_{x\in\mathbb{R}}\left\{(t_{0}+\delta)\wedge x~:~\Lambda(x)\geqslant 1-\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}},~\mathbb{Q}\mbox{-almost surely}\right\}$
$=\lim_{\delta\downarrow 0}\sup_{x\in\mathbb{R}}\left\{t_{0}\wedge x~:~\Lambda(x)\geqslant 1-\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}},~\mathbb{Q}\mbox{-almost surely}\right\}=R(t_{0},\mathbb{Q}).$
Monotonicity is straightforward, and quasi-concavity is implied by monotonicity. Statement (iii) is clear because $\inf_{t\in\mathbb{R}}R(t,\mathbb{Q})=$-$\infty$ for all $\mathbb{Q}\in\mathcal{M}_{1,f}$.
(iv) For all $t_{1}\geqslant t_{2}$ and $\mathbb{Q}\in\mathcal{M}_{1,f}$,
$R(t_{1},\mathbb{Q})-R(t_{2},\mathbb{Q})=$ $~t_{1}\wedge\sup_{x\in\mathbb{R}}\left\{x~:~\Lambda(x)\geqslant 1-\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}},~\mathbb{Q}\mbox{-almost surely}\right\}$
$-t_{2}\wedge\sup_{x\in\mathbb{R}}\left\{x~:~\Lambda(x)\geqslant 1-\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}},~\mathbb{Q}\mbox{-almost surely}\right\}\leqslant t_{1}-t_{2}.$
The proof is complete. ∎
Remark 2.
Suppose that $\Lambda:\mathbb{R}\to[0,1]$ is decreasing and left-continuous. The function $R$ we obtained in (18) is a special case of that obtained by Theorem 3.1 of Cerreia-Vioglio et al. (2011) for quasi-convex cash-subadditive risk measures:
$R(t,\mathbb{Q})=\inf\{\mathrm{ES}_{\Lambda}(Y):\mathbb{E}_{\mathbb{Q}}[Y]=t\},~(t,\mathbb{Q})\in\mathbb{R}\times\mathcal{M}_{1,f}.$ (19)
Theorem 3 automatically implies (19). Below, we show another self-contained proof for (19) to provide more mathematical insight. This proof can be seen as an alternative proof for Theorem 3 with $\Lambda$ being left-continuous. For any $Y\in L^{\infty}$, due to boundedness of $Y$, there exist $a,b\in\mathbb{R}$ with $a<b$, such that
$\mathrm{ES}_{\Lambda}(Y)=\sup_{y\in[a,b]}\left(\mathrm{ES}_{\Lambda(y)}(Y)\wedge y\right).$ (20)
For any $y_{1},y_{2}\in[a,b]$ with $y_{1}\leqslant y_{2}$ and $\gamma\in[0,1]$, because $y\mapsto\mathrm{ES}_{\Lambda(y)}$ is decreasing, we have
$\mathrm{ES}_{\Lambda(\gamma y_{1}+(1-\gamma)y_{2})}(Y)\wedge\left(\gamma y_{1}+(1-\gamma)y_{2}\right)$
$=\left\{\begin{array}[]{ll}\mathrm{ES}_{\Lambda(\gamma y_{1}+(1-\gamma)y_{2})}(Y)\geqslant\mathrm{ES}_{\Lambda(y_{2})}(Y),&\mbox{if}~~\mathrm{ES}_{\Lambda(\gamma y_{1}+(1-\gamma)y_{2})}(Y)\leqslant\gamma y_{1}+(1-\gamma)y_{2},\\ \gamma y_{1}+(1-\gamma)y_{2}\geqslant y_{1},&\mbox{otherwise}\end{array}\right$.
$\geqslant\left(\mathrm{ES}_{\Lambda(y_{1})}(Y)\wedge y_{1}\right)\wedge\left(\mathrm{ES}_{\Lambda(y_{2})}(Y)\wedge y_{2}\right).$
Thus the function $y\mapsto\mathrm{ES}_{\Lambda(y)}(Y)\wedge y$ is quasi-concave. For any $(t,\mathbb{Q})\in\mathbb{R}\times\mathcal{M}_{1,f}$, it is clear that the set $\{Y\in L^{\infty}:\mathbb{E}_{\mathbb{Q}}[Y]=t\}$ is convex and the mapping $Y\mapsto\mathrm{ES}_{\Lambda(y)}(Y)\wedge y$ is convex due to convexity of $\mathrm{ES}$. Hence,
$\inf\left\{\mathrm{ES}_{\Lambda}(Y):\mathbb{E}_{\mathbb{Q}}[Y]=t\right\}$
$=\inf\left\{\sup_{y\in[a,b]}\left(\mathrm{ES}_{\Lambda(y)}(Y)\wedge y\right):\mathbb{E}_{\mathbb{Q}}[Y]=t\right\}$ [by (20)]
$=\sup_{y\in[a,b]}\inf_{\mathbb{E}_{\mathbb{Q}}[Y]=t}\left(\mathrm{ES}_{\Lambda(y)}(Y)\wedge y\right)$ [minimax theorem (Fan, 1953)]
$=\sup_{y\in[a,b]}\left(\inf_{\mathbb{E}_{\mathbb{Q}}[Y]=t}\mathrm{ES}_{\Lambda(y)}(Y)\wedge y\right)$ [exchangeability of two infimums]
$=\sup_{y\in[a,b]}\left(\inf_{c\in\mathbb{R}}\inf_{Y\in L^{\infty}}\left(\mathrm{ES}_{\Lambda(y)}(Y)-c(\mathbb{E}_{\mathbb{Q}}[Y]-t)\right)\wedge y\right)$ [Lagrange duality]
$=\sup_{y\in[a,b]}\left(\inf_{c\in\mathbb{R}}\left(ct-\alpha(c\mathbb{Q})\right)\wedge y\right)$ $\left[\displaystyle\alpha(\mathbb{Q})=\sup_{Y\in L^{\infty}}(\mathbb{E}_{\mathbb{Q}}[Y]-\mathrm{ES}_{\Lambda(y)}(Y))\right]$
$=\sup_{y\in[a,b]}\left(\left(t-\alpha(\mathbb{Q})\right)\wedge y\right).$ [cash additivity of $\mathrm{ES}$]
For all $\mathbb{Q}\in\mathcal{M}_{1,f}$ and $y\in[a,b]$, by Corollary 4.19 and Theorem 4.52 of Föllmer and Schied (2016), we have $\alpha(\mathbb{Q})=0$ if $\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}\leqslant 1/(1-\Lambda(y))$, $\mathbb{P}$-almost surely and $\alpha(\mathbb{Q})=\infty$ otherwise. Therefore, we have
$\sup_{y\in[a,b]}\left(\left(t-\alpha(\mathbb{Q})\right)\wedge y\right)=\max_{x\in\mathbb{R}}\left\{t\wedge x~:~\Lambda(x)\geqslant 1-\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}},~\mathbb{Q}\mbox{-almost surely}\right\}.$
5 The Rockafellar–Uryasev formula and optimization
5.1 Representing Lambda ES as a minimization
The well-known relation between VaR and ES obtained by Rockafellar and Uryasev (2002) as shown in (3) provides a promising solution to ES-based optimization problems. Let $I(\overline{\mathbb{R}})$ be the set of all closed real intervals, including intervals of $\ell$$\infty$], $[\ell,\infty)$, $\infty$$\ell$], $(-\infty,\ell]$, and $[\ell,\ell]=\{\ell\}$ for $\ell\in\overline{\mathbb{R}}$. A pair of risk measures $(\phi,\rho):\mathcal{X}\to I(\overline{\mathbb{R}})\times\overline{\mathbb{R}}$ is called a Bayes pair (Embrechts et al., 2021) if for some loss function $S:\overline{\mathbb{R}}^{2}\to\overline{\mathbb{R}}$,
$\phi(X)=\operatorname*{arg\,min}_{a\in\overline{\mathbb{R}}}\mathbb{E}[S(a,X)],~\mbox{and}~\rho(X)=\min_{a\in\overline{\mathbb{R}}}\mathbb{E}[S(a,X)],~~X\in\mathcal{X}.$
If $\phi$ is further cash additive (naturally defined for interval-valued functions), then we call $\rho$ a Bayes risk measure, and $\phi$ the corresponding Bayes estimator. It is clear that $(\mathrm{VaR},\mathrm{ES})$ is a Bayes pair by (3). A natural question is whether we can also write $\mathrm{ES}_{\Lambda}$ as the minimum of some loss function and find its corresponding minimizer. Ideally, we may expect to find the relation between $\mathrm{VaR}_{\Lambda}$ and $\mathrm{ES}_{\Lambda}$ similar to $(\mathrm{VaR},\mathrm{ES})$ in optimization. Below we show a representation of $\mathrm{ES}_{\Lambda}$ based on the relation (3), which we call the RU representation of $\mathrm{ES}_{\Lambda}$ for simplicity. Define the mapping $T_{\Lambda}:\mathbb{R}\times\mathbb{R}\times L^{\infty}\to(-\infty,\infty]$ as
$T_{\Lambda}:(a,x,X)\mapsto\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(X-a)_{+}\right]\vee x.$ (21)
In the next result, we also study convexity of $T_{\Lambda}$ with respect to different variables. We use the term “joint convexity” when there is more than one variable to emphasize that the property is different from convexity in each variable.
Theorem 4.
Let $\Lambda:\mathbb{R}\to[0,1]$ be a right-continuous decreasing function and $T_{\Lambda}$ be given in (21). We have
$\mathrm{ES}_{\Lambda}(X)=\min_{(a,x)\in\mathbb{R}^{2}}T_{\Lambda}(a,x,X)=\min_{(a,x)\in\mathbb{R}^{2}}\left\{\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(X-a)_{+}\right]\vee x\right\},~~X\in L^{\infty},$ (22)
where the minima are obtained at $x^{*}=\mathrm{ES}_{\Lambda}(X)$ and
$a^{*}\left\{\begin{array}[]{ll}\in[\mathrm{VaR}_{\Lambda(x^{*})}(X),\mathrm{VaR}^{+}_{\Lambda(x^{*})}(X)],&\mbox{if}~\Lambda(x^{*})\in[0,1),\\ =\mathrm{VaR}_{1}(X),&\mbox{if}~\Lambda(x^{*})=1.\end{array}\right$.
Moreover,
  1. (i)
    $T_{\Lambda}(a,x,X)$ is jointly convex in $(a,X)\in\mathbb{R}\times L^{\infty}$ for all $x\in\mathbb{R}$;
  2. (ii)
    $T_{\Lambda}(a,x,X)$ is convex in $x\in\mathbb{R}$ for all $(a,X)\in\mathbb{R}\times L^{\infty}$ if and only if the function $x\mapsto 1/(1-\Lambda(x))$ is convex;
  3. (iii)
    the following statements are equivalent:
    1. (a)
      $T_{\Lambda}(a,x,X)$ is jointly convex in $(a,x)\in\mathbb{R}^{2}$ for all $X\in L^{\infty}$;
    2. (b)
      $T_{\Lambda}(a,x,X)$ is jointly quasi-convex in $(a,x)\in\mathbb{R}^{2}$ for all $X\in L^{\infty}$;
    3. (c)
      $T_{\Lambda}(a,x,X)$ is jointly convex in $(a,x,X)\in\mathbb{R}\times\mathbb{R}\times L^{\infty}$;
    4. (d)
      $T_{\Lambda}(a,x,X)$ is jointly quasi-convex in $(a,x,X)\in\mathbb{R}\times\mathbb{R}\times L^{\infty}$;
    5. (e)
      $\Lambda$ is constant on $\mathbb{R}$.
Proof.
For any $X\in L^{\infty}$, we have by Theorem 2, formulation (3), and Remark 1 that
$\mathrm{ES}_{\Lambda}(X)=\min_{x\in\mathbb{R}}\left\{\mathrm{ES}_{\Lambda(x)}(X)\vee x\right\}$ $=\min_{x\in\mathbb{R}}\left\{\min_{a\in\mathbb{R}}\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(X-a)_{+}\right]\vee x\right\}$
$=\min_{x\in\mathbb{R}}\min_{a\in\mathbb{R}}\left\{\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(X-a)_{+}\right]\vee x\right\}.$
For the optimization problem above, the minimizer $x^{*}=\mathrm{ES}_{\Lambda}(X)$ is obtained by definition (12), and the minimizer $a^{*}$ is obtained by (6).
Statement (i) is straightforward. We prove statements (ii) - (iv).
(ii) The “if” part is clear by the convexity of $x\mapsto 1/(1-\Lambda(x))$. To show the “only if” part, suppose (21) is convex in $x$. Let
$\bar{x}=\inf\{x\in\mathbb{R}:\Lambda(x)=0\}.$
Right-continuity of $\Lambda$ yields that $\Lambda(\bar{x})=0$. We first prove $x\mapsto 1/(1-\Lambda(x))$ is convex in $x\in(-\infty,\bar{x})$. Suppose for contradiction that
$\frac{1}{1-\Lambda\left(\frac{x_{0}+y_{0}}{2}\right)}>\frac{1}{2(1-\Lambda(x_{0}))}+\frac{1}{2(1-\Lambda(y_{0}))},~~\mbox{for some}~x_{0},y_{0}\in(-\infty,\bar{x}).$ (23)
Because it is clear that
$\lim_{a\downarrow-\infty}\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(X-a)_{+}\right]=\infty,~~\mbox{for all}~x\in(-\infty,\bar{x}),$ (24)
there exists $a_{0}\in\mathbb{R}$, such that
$\mathbb{E}\left[a_{0}+\frac{1}{1-\Lambda(x_{0})}(X-a_{0})_{+}\right]\geqslant x_{0},~\mathbb{E}\left[a_{0}+\frac{1}{1-\Lambda(y_{0})}(X-a_{0})_{+}\right]\geqslant y_{0},$ (25)
and $\mathbb{E}\left[a_{0}+\frac{1}{1-\Lambda\left((x_{0}+y_{0})/2\right)}(X-a_{0})_{+}\right]\geqslant\frac{x_{0}+y_{0}}{2}.$
(23) and (25) together contradict the fact that (21) is convex in $x$. Therefore, the function $x\mapsto 1/(1-\Lambda(x))$ is convex in $x\in(-\infty,\bar{x})$.
Next, we show that $x\mapsto 1/(1-\Lambda(x))$ is convex in $x\in\mathbb{R}$. Because $x\mapsto 1/(1-\Lambda(x))$ is decreasing, it suffices to show that $x\mapsto 1/(1-\Lambda(x))$ is continuous at $\bar{x}$ if $\bar{x}<\infty$. Suppose for contradiction that $\Lambda(\bar{x}-)>0$. Because of (24), there exists $a_{1}\in(-\infty,\mathrm{ess\mbox{-}sup}(X))$, such that
$\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda(\bar{x}-)}(X-a_{1})_{+}\right]>\bar{x}.$
It is clear that $\mathbb{E}[(X-a_{1})_{+}]>0$ and thus
$\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda(\bar{x}-)}(X-a_{1})_{+}\right]>\mathbb{E}\left[a_{1}+(X-a_{1})_{+}\right].$
It follows that
$\lim_{x\uparrow\bar{x}}\frac{\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda(x)}(X-a_{1})_{+}\right]\vee x}{2}+\frac{\mathbb{E}\left[a_{1}+(X-a_{1})_{+}\right]\vee\bar{x}}{2}$
$=\frac{\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda(\bar{x}-)}(X-a_{1})_{+}\right]}{2}+\frac{\mathbb{E}\left[a_{1}+(X-a_{1})_{+}\right]\vee\bar{x}}{2}$
$<\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda(\bar{x}-)}(X-a_{1})_{+}\right]$
$=\lim_{x\uparrow\bar{x}}\left\{\mathbb{E}\left[a_{1}+\frac{1}{1-\Lambda((x+\bar{x})/2)}(X-a_{1})_{+}\right]\vee\frac{x+\bar{x}}{2}\right\}.$
This contradicts with the convexity of (21). Therefore, $x\mapsto 1/(1-\Lambda(x))$ is convex in $x\in\mathbb{R}$.
Next, we prove statement (iii). It is straightforward that (a) $\Rightarrow$ (b) and (c) $\Rightarrow$ (d).
“(e) $\Rightarrow$ (a)”: This follows by the convexity of (21) in $a\in\mathbb{R}$ and the fact that an increasing convex transform of a convex function is still convex. We show the “only if” part.
“(b) $\Rightarrow$ (e)”: Suppose that (21) is jointly quasi-convex in $(a,x)\in\mathbb{R}^{2}$ for all $X\in L^{\infty}$. Suppose for contradiction that $\Lambda$ is decreasing and non-constant on $\mathbb{R}$. There exists $x,y,t\in\mathbb{R}$ with $y<x\leqslant t$, such that $\Lambda(y)\geqslant\Lambda((x+y)/2)>\Lambda(x)$. Take $a<b=t$, and $X\in L^{\infty}$ with $\mathbb{P}(X=a)=1-\mathbb{P}(X=t)=\Lambda(x)$. Because $\Lambda(x)<1$, we have
$\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(X-a)_{+}\right]\vee x=\mathbb{E}\left[b+\frac{1}{1-\Lambda(y)}(X-b)_{+}\right]\vee y=t,$
whereas
$\mathbb{E}\left[\frac{a+b}{2}+\frac{1}{1-\Lambda((x+y)/2)}\left(X-\frac{a+b}{2}\right)_{+}\right]\vee\frac{x+y}{2}$
$=\left\{\frac{a+t}{2}+\frac{1-\Lambda(x)}{1-\Lambda((x+y)/2)}\left(t-\frac{a+t}{2}\right)\right\}\vee\frac{x+y}{2}>t.$
This contradicts the joint quasi-convexity of (21) in $(a,x)\in\mathbb{R}^{2}$, and thus $\Lambda$ is constant on $\mathbb{R}$.
“(e) $\Rightarrow$ (c)”: This follows by statement (i) and the fact that an increasing convex transform of a convex function is still convex.
“(d) $\Rightarrow$ (e)”: This follows directly by the proof for the implication “(b) $\Rightarrow$ (e)”. ∎
For the RU representation of $\Lambda$-ES, Theorem 4 indicates that we do not guarantee joint convexity (or quasi-convexity) of the objective (21) in $(a,x)\in\mathbb{R}^{2}$ or $(a,x,X)\in\mathbb{R}\times\mathbb{R}\times L^{\infty}$ unless $\Lambda$ is a constant (i.e., $\mathrm{ES}_{\Lambda}$ is an ES). Nevertheless, the mapping (21) is convex in $x\in\mathbb{R}$ when the function $x\mapsto 1/(1-\Lambda(x))$ is convex. We give two examples satisfying this condition: $\Lambda$ is a constant (corresponding to the usual ES) and $\Lambda(x)=(e^{ax}+1)^{-1}$ for a>0a>0.444The example $\Lambda(x)=(e^{ax}+1)^{-1}$, $a>0$, $x\in\mathbb{R}$, provides a suggestion of a non-trivial choice of the $\mathrm{ES}_{\Lambda}$ risk measure to use in practice. By Theorem 4, an obvious advantage of such a choice is that it makes the objective of a practical $\Lambda$-ES-based optimization problem convex in the variable $x$. Moreover, Theorem 4 shows that $\mathrm{ES}_{\Lambda}$ has a similar feature to a Bayes risk measure, as it can be represented as the minimum of an expected loss function with an additional transformation. The corresponding minimizer is the interval of the left- and right-quantiles at the level of $\Lambda(x^{*})$ instead of $\mathrm{VaR}_{\Lambda},\mathrm{VaR}^{+}_{\Lambda}$].
A possible direction to explore the issue of Bayes pair is through the scoring function of $\Lambda$-$\mathrm{VaR}$. [^5] A set-valued functional $\rho:\mathcal{X}\to 2^{\mathbb{R}}$ is called elicitable on $\mathcal{X}$ if there exists a function (scoring function) $S:\overline{\mathbb{R}}^{2}\to\overline{\mathbb{R}}$, such that $\rho(X)=\operatorname*{arg\,min}_{a\in\overline{\mathbb{R}}}\mathbb{E}[S(a,X)],~~X\in\mathcal{X}.$ Bellini and Bignozzi (2015) and Burzoni et al. (2017) showed that for a decreasing function $\Lambda:\mathbb{R}\to(0,1)$, $\mathrm{VaR}_{\Lambda}$ is elicitable with the scoring function
$S_{\Lambda}(a,y)=(a-y)_{+}-\int_{y}^{a}\Lambda(t)\,\mathrm{d}t=(y-a)_{+}-\int_{a}^{y}(1-\Lambda(t))\,\mathrm{d}t,~~a\in\overline{\mathbb{R}},~y\in\mathbb{R}.$
The pair of risk measures we get with the above scoring function is $(\mathrm{VaR}_{\Lambda},\rho)$, where
$\rho_{\Lambda}(X)=\min_{a\in\overline{\mathbb{R}}}\mathbb{E}[cS_{\Lambda}(a,X)+f(X)],~~X\in L^{\infty},$
for some constant $c>0$ and real function $f:\mathbb{R}\to\mathbb{R}$. The risk measure $(\mathrm{VaR}_{\Lambda},\rho_{\Lambda})$ is not a Bayes pair because $\mathrm{VaR}_{\Lambda}$ is not cash additive in general. Moreover, we find that $\rho_{\Lambda}$ cannot satisfy quasi-convexity, normalization, and $\rho_{\Lambda}\geqslant\mathrm{VaR}_{\Lambda}$ simultaneously, and thus does not coincide with $\mathrm{ES}_{\Lambda}$ for any choices of $c$ and $f$. We put the detailed arguments for this conflict in Appendix A.
5.2 Optimization with Lambda ES
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function. Based on the theoretical results in the previous sections, we demonstrate general ideas of solving optimization problems with $\Lambda$-ES as a constraint or an objective. For $n\in\mathbb{N}$, let $\mathbf{L}=(L_{1},\dots,L_{n})\in(L^{\infty})^{n}$ represent a vector of losses, $\bm{\theta}\in\Theta$ represent a decision variable, where $\Theta$ is a convex set of actions, and $f:\Theta\times\mathbb{R}^{n}\to\mathbb{R}$ represent a loss function. For a typical example in finance, we can use $\mathbf{L}$ as the vector of losses from multiple assets and $\bm{\theta}$ as a portfolio weight vector.
First, we are interested in a problem where the decision maker aims at minimizing an objective risk measure $\rho:L^{\infty}\to\overline{\mathbb{R}}$ of the aggregate loss $f(\bm{\theta},\mathbf{L})$, guaranteeing that the $\Lambda$-ES of the total loss does not exceed a pre-specified value $\ell\in\mathbb{R}$. Namely, we consider the following optimization problem:
$\min_{\bm{\theta}\in\Theta}\rho(f(\bm{\theta},\mathbf{L}))~~~\mbox{subject to}~\mathrm{ES}_{\Lambda}(f(\bm{\theta},\mathbf{L}))\leqslant\ell.$ (26)
The following result provides a possible direction to simplify the problem above.
Proposition 4.
Let $\Lambda:\mathbb{R}\to[0,1]$ be a right-continuous decreasing function. The constraint $\mathrm{ES}_{\Lambda}(f(\bm{\theta},\mathbf{L}))\leqslant\ell$ in (26) is equivalent to $\mathrm{ES}_{\Lambda(\ell)}(f(\bm{\theta},\mathbf{L}))\leqslant\ell$.
Proof.
By definition, we have
$\mathrm{ES}_{\Lambda}(f(\bm{\theta},\mathbf{L}))\leqslant\ell$ $\iff\sup_{x\in\mathbb{R}}\left(\mathrm{ES}_{\Lambda(x)}(f(\bm{\theta},\mathbf{L}))\wedge x\right)\leqslant\ell$
$\iff\mbox{for all }x\in\mathbb{R}:~\mathrm{ES}_{\Lambda(x)}(f(\bm{\theta},\mathbf{L}))\wedge x\leqslant\ell$
$\iff\mbox{for all }x\in\mathbb{R}:~\mathrm{ES}_{\Lambda(x)}(f(\bm{\theta},\mathbf{L}))\leqslant\ell~\mbox{or}~x\leqslant\ell$
$\iff\sup_{x>\ell}\mathrm{ES}_{\Lambda(x)}(f(\bm{\theta},\mathbf{L}))\leqslant\ell\iff\mathrm{ES}_{\Lambda(\ell)}(f(\bm{\theta},\mathbf{L}))\leqslant\ell,$
where the last equivalence holds by right-continuity of $\Lambda$. ∎
Proposition 4 implies that an optimization problem with a $\Lambda$-ES constraint below level $\ell\in\mathbb{R}$ can be equivalently converted to a problem with the constraint on $\mathrm{ES}_{\Lambda(\ell)}$ below the same level $\ell$. ES-constrained optimization problem has been studed extensively in the literature (see e.g., Krokhmal et al., 2002).
Another natural question to study is how to minimize $\Lambda$-ES as an objective. We consider the following optimization problem:
$\min_{\bm{\theta}\in\Theta}\mathrm{ES}_{\Lambda}(f(\bm{\theta},\mathbf{L})).$ (27)
In the current work, we aim to provide general insights into solving $\Lambda$-ES-based optimization problems, as stated in the result below. In specific problems, we may also consider some constraints along with the problem (27), which does not change the nature of our equivalence result. Define the mapping $\widetilde{T}_{\Lambda}:\Theta\times\mathbb{R}\times\mathbb{R}\times(L^{\infty})^{n}\to\overline{\mathbb{R}}$ as
$\widetilde{T}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})=\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(f(\bm{\theta},\mathbf{L})-a)_{+}\right]\vee x.$ (28)
We have the following result.
Proposition 5.
For a right-continuous decreasing function $\Lambda:\mathbb{R}\to[0,1]$,
$\min_{\bm{\theta}\in\Theta}\mathrm{ES}_{\Lambda}(f(\bm{\theta},\mathbf{L}))=\min_{(\bm{\theta},a,x)\in\Theta\times\mathbb{R}\times\mathbb{R}}\left\{\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(f(\bm{\theta},\mathbf{L})-a)_{+}\right]\vee x\right\}.$ (29)
Moreover, for $\widetilde{T}_{\Lambda}$ defined in (28),
  1. (i)
    $\widetilde{T}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})$ is convex in $a\in\mathbb{R}$ for all $(\bm{\theta},x,\mathbf{L})\in\Theta\times\mathbb{R}\times(L^{\infty})^{n}$.
  2. (ii)
    if in addition, $\Lambda$ is not constantly $1$, then $\widetilde{T}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})$ is jointly convex in $(\bm{\theta},\mathbf{L})\in\Theta\times(L^{\infty})^{n}$ for all $(a,x)\in\mathbb{R}^{2}$ if and only if $f(\bm{\theta},\mathbf{L})$ is jointly convex in $(\bm{\theta},\mathbf{L})$.
  3. (iii)
    $\widetilde{T}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})$ is convex in $x\in\mathbb{R}$ for all $(\bm{\theta},a,\mathbf{L})\in\Theta\times\mathbb{R}\times(L^{\infty})^{n}$ if and only if the function $x\mapsto 1/(1-\Lambda(x))$ is convex.
Proof.
The equation (29) holds directly by Theorem 4. Statements (i) and (iii) follow by Theorem 4. Statement (ii) is straightforward. ∎
6 Extensions to the space of integrable random variabls
In this section, we extend our discussions on ES and $\Lambda$-ES from the space of $L^{\infty}$ to $L^{1}$. Similarly to the corresponding definitions on $L^{\infty}$, we define ES at level $\alpha\in[0,1]$ as the mapping $\mathrm{ES}_{\alpha}:L^{1}\to\overline{\mathbb{R}}$ given by
$\mathrm{ES}_{\alpha}(X)=\frac{1}{1-\alpha}\int_{\alpha}^{1}\mathrm{VaR}_{\beta}(X)\mathrm{d}\beta\mbox{ for $\alpha\in[0,1)$},$
and $\mathrm{ES}_{1}(X)=\mathrm{VaR}_{1}(X)$; for a decreasing function $\Lambda:\mathbb{R}\to[0,1]$, we define $\mathrm{ES}_{\Lambda}:L^{1}\to\overline{\mathbb{R}}$ as
$\mathrm{ES}_{\Lambda}(X)=\sup_{x\in\mathbb{R}}\left\{\mathrm{ES}_{\Lambda(x)}(X)\wedge x\right\},~~X\in L^{1}.$
Some of the results in the previous sections can be naturally extended to $L^{1}$, whereas others only hold under a weakened setup, for which we provide independent proofs for completeness. For the convenience of our discussion, we first note that the properties in Propositions 2 and 3 still hold for $\mathrm{ES}_{\Lambda}$ on $L^{1}$ by the same arguments in its proof replacing $L^{\infty}$ by $L^{1}$.
6.1 Finiteness of Lambda ES
Below we show the finiteness of $\Lambda$-ES on $L^{1}$. As a result, the risk measure $\mathrm{ES}_{\Lambda}$ is always well-defined (possibly being $\infty$) on $L^{1}$.
Proposition 6.
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function. The mapping $\mathrm{ES}_{\Lambda}:L^{1}\to\overline{\mathbb{R}}$ satisfies
$\infty<\mathbb{E}[X]\leqslant\mathrm{ES}_{\Lambda}(X)\leqslant\mathrm{ES}_{1}(X),~~~X\in L^{1}.$
In particular, $\mathrm{ES}_{\Lambda}(X)$ is finite on $L^{1}$ if and only if $\mathrm{VaR}_{1}(X)<\infty$ or $\Lambda$ is not constantly $1$.
Proof.
The relation $\infty<\mathbb{E}\leqslant\mathrm{ES}_{\Lambda}\leqslant\mathrm{ES}_{1}$ holds by item (a) of Proposition 2 on $L^{1}$. We prove the “if” part of the second statement, whose “only if” part is straightforward.
For any $X\in L^{1}$, first, suppose that $\mathrm{VaR}_{1}(X)<\infty$. It is straightforward that $\mathrm{ES}_{\Lambda}(X)\leqslant\mathrm{ES}_{1}(X)=\mathrm{VaR}_{1}(X)<\infty$. Next, suppose that $\Lambda$ is not constantly $1$. There exists $x_{0}\in\mathbb{R}$, such that $0\leqslant\Lambda(x_{0})<1$. It follows that $\mathrm{ES}_{\Lambda(x_{0})}(X)<\infty$. By (13),
$\mathrm{ES}_{\Lambda}(X)=\inf_{x\in\mathbb{R}}\left(\mathrm{ES}_{\Lambda(x)}(X)\vee x\right)\leqslant\mathrm{ES}_{\Lambda(x_{0})}(X)\vee x_{0}<\infty.$
The proof is complete. ∎
6.2 Dominance of ES and Lambda ES
Here, we examine the $L^{1}$ versions of the dominance results in Theorems 1 and 2. Theorem 1 does not hold in general if we extend the space of $\mathrm{ES}_{\alpha}$ from $L^{\infty}$ to $L^{1}$ because the dominance may fail at $\alpha=0$. A counterexample can be: Let $\rho:L^{1}\to\overline{\mathbb{R}}$ be a risk measure defined as follows.
$\rho(X)=\left\{\begin{array}[]{ll}\mathrm{ess\mbox{-}sup}(X),&\mbox{if }\mathrm{ess\mbox{-}sup}(X)=\infty~\mbox{or}~\mathrm{ess\mbox{-}inf}(X)>-\infty,\\ -\infty,&\mbox{if }\mathrm{ess\mbox{-}sup}(X)<\infty~\mbox{and}~\mathrm{ess\mbox{-}inf}(X)=-\infty,\end{array}\right.~X\in L^{1}.$
One can check that $\rho$ is quasi-convex, law-invariant, and $\rho\geqslant\mathrm{VaR}^{+}_{0}$. However, the condition $\rho\geqslant\mathbb{E}$ fails. Therefore, $\mathbb{E}$ is not the smallest quasi-convex and law-invariant risk measure dominating $\mathrm{VaR}^{+}_{0}$ and thus (8) in Theorem 1 fails for $\alpha=0$.
Below, we state the dominance results for ES over VaR and $\Lambda$-ES over $\Lambda$-VaR on the space of $L^{1}$. Both results rely on slightly stronger assumptions than Theorems 1 and 2 regarding the case of $\alpha=0$. Write $\underline{L}^{1}$ as the set of all random variables in $L^{1}$ that are essentially bounded from below.
Theorem 5.
For any $\alpha\in(0,1]$,
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{1}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (30)
For any $\alpha\in(0,1)$,
$\mathrm{ES}_{\alpha}=\min\{\rho:L^{1}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}^{+}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (31)
For any $\alpha\in[0,1)$,
$\mathrm{ES}_{\alpha}=\min\{\rho:\underline{L}^{1}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}^{+}_{\alpha}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (32)
Proof.
The proofs for (30), (31), and the case of $\alpha\in(0,1)$ for (32) follow directly from those for Theorem 1 whose arguments still hold on $L^{1}$. We only need to prove (32) for $\alpha=0$.
For any $n\in\mathbb{N}$ and $X\in\underline{L}^{1}$ with distribution $F$, write $K_{n}=\sqrt{n}+\mathrm{VaR}^{+}_{0}(X)$. By Corollary A.3 of Embrechts et al. (2015), there exist $\widetilde{X}_{i}\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}F$, $i\in[n]$, such that
$\left|\frac{1}{n}\sum^{n}_{i=1}\left(\widetilde{X}_{i}\wedge K_{n}\right)-\mathbb{E}\left[X\wedge K_{n}\right]\right|\leqslant\frac{\sqrt{n}}{n}=\frac{1}{\sqrt{n}}.$
It follows that
$\frac{1}{n}\sum^{n}_{i=1}\left(\widetilde{X}_{i}\wedge K_{n}\right)\geqslant\mathbb{E}\left[X\wedge K_{n}\right]-\frac{1}{\sqrt{n}},$
which implies that
$\mathbb{E}\left[X\wedge K_{n}\right]\geqslant\frac{1}{n}\mathrm{VaR}^{+}_{0}\left(\left(\widetilde{X}_{1}\wedge K_{n}\right)+\cdots+\left(\widetilde{X}_{n}\wedge K_{n}\right)\right)\geqslant\mathbb{E}\left[X\wedge K_{n}\right]-\frac{1}{\sqrt{n}}.$
Letting $n\to\infty$, by monotone convergence theorem, we have
$\frac{1}{n}\sup\left\{\mathrm{VaR}^{+}_{0}\left({X}_{1}+\cdots+{X}_{n}\right):X_{i}\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}F,~i\in[n]\right\}\to\mathbb{E}[X].$
The rest of the proof follows from that for Theorem 1. ∎
The proof of the following result follows the same arguments as that for Theorem 2, extending the space from $L^{\infty}$ to $L^{1}$ or $\underline{L}^{1}$.
Theorem 6.
The following statements hold.
  1. (i)
    For a decreasing function $\Lambda:\mathbb{R}\to(0,1]$, the smallest quasi-convex and law-invariant risk measure on $L^{1}$ dominating $\mathrm{VaR}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^1Rρ⩾VaR_Λ and $\rho$ is quasi-convex and law-invariant}.
  2. (ii)
    For a decreasing function $\Lambda:\mathbb{R}\to(0,1)$, the smallest quasi-convex and law-invariant risk measure on $L^{1}$ dominating $\mathrm{VaR}^{+}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^1Rρ⩾VaR^+_Λ and $\rho$ is quasi-convex and law-invariant}.
  3. (iii)
    For a decreasing function $\Lambda:\mathbb{R}\to[0,1)$, the smallest quasi-convex and law-invariant risk measure on $\underline{L}^{1}$ dominating $\mathrm{VaR}^{+}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^1Rρ⩾VaR^+_Λ and $\rho$ is quasi-convex and law-invariant}.
Moreover, the identity holds for all decreasing functions $\Lambda:\mathbb{R}\to[0,1]$:
$\mathrm{ES}_{\Lambda}(X)=\inf_{x\in\mathbb{R}}\left(\mathrm{ES}_{\Lambda(x)}(X)\vee x\right),~~~X\in L^{1}.$
6.3 Dual and RU representations for Lambda ES
In this section, we restate the dual representation (Theorem 3) and the RU representation (Theorem 4) for $\Lambda$-ES on $L^{1}$. The proofs of both results below follow from the same arguments as those for Theorems 3 and 4 by replacing $L^{\infty}$ with $L^{1}$.
Theorem 7.
For any decreasing function $\Lambda:\mathbb{R}\to[0,1]$, the risk measure $\mathrm{ES}_{\Lambda}$ adopts the following representation:
$\mathrm{ES}_{\Lambda}(X)=\sup_{\mathbb{Q}\in\mathcal{M}_{1,f}}R(\mathbb{E}_{\mathbb{Q}}[X],\mathbb{Q}),~~~X\in L^{1},$ (33)
where for $(t,\mathbb{Q})\in\mathbb{R}\times\mathcal{M}_{1,f}$,
$R(t,\mathbb{Q})=\sup_{x\in\mathbb{R}}\left\{t\wedge x~:~\Lambda(x)\geqslant 1-\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}},~\mathbb{Q}\mbox{-almost surely}\right\}.$
Moreover, the following are true:
  1. (i)
    The supremum in (33) can be changed to a maximum if $\Lambda$ is left-continuous.
  2. (ii)
    $(t,\mathbb{Q})\mapsto R(t,\mathbb{Q})$ is upper semicontinuous, quasi-concave, and increasing in $t$;
  3. (iii)
    $\inf_{t\in\mathbb{R}}R(t,\mathbb{Q})=\inf_{t\in\mathbb{R}}R(t,\mathbb{Q}^{\prime})$ for all $\mathbb{Q},\mathbb{Q}^{\prime}\in\mathcal{M}_{1,f}$;
  4. (iv)
    $R(t_{1},\mathbb{Q})-R(t_{2},\mathbb{Q})\leqslant t_{1}-t_{2}$ for all $t_{1}\geqslant t_{2}$ and $\mathbb{Q}\in\mathcal{M}_{1,f}$.
With a slight abuse of notation, we define the mapping $T_{\Lambda}:\overline{\mathbb{R}}\times\mathbb{R}\times L^{1}\to\overline{\mathbb{R}}$ as
$T_{\Lambda}:(a,x,X)\mapsto\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(X-a)_{+}\right]\vee x.$ (34)
Proposition 7.
For any right-continuous decreasing function $\Lambda:\mathbb{R}\to[0,1]$, the risk measure $\mathrm{ES}_{\Lambda}$ can be represented as follows.
$\mathrm{ES}_{\Lambda}(X)=\min_{(a,x)\in\overline{\mathbb{R}}\times\mathbb{R}}T(a,x,X)=\min_{(a,x)\in\overline{\mathbb{R}}\times\mathbb{R}}\left\{\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(X-a)_{+}\right]\vee x\right\},~~X\in L^{1},$
where the minima are obtained at $x^{*}=\mathrm{ES}_{\Lambda}(X)$ and
$a^{*}\left\{\begin{array}[]{ll}\in[\mathrm{VaR}_{\Lambda(x^{*})}(X),\mathrm{VaR}^{+}_{\Lambda(x^{*})}(X)],&\mbox{if}~\Lambda(x^{*})\in[0,1),\\ =\mathrm{VaR}_{1}(X),&\mbox{if}~\Lambda(x^{*})=1.\end{array}\right$.
Moreover, for $T_{\Lambda}$ defined in (34),
  1. (i)
    $T_{\Lambda}(a,x,X)$ is jointly convex in $(a,X)\in\overline{\mathbb{R}}\times L^{1}$ for all $x\in\mathbb{R}$;
  2. (ii)
    $T_{\Lambda}(a,x,X)$ is convex in $x\in\mathbb{R}$ for all $(a,X)\in\overline{\mathbb{R}}\times L^{1}$ if and only if the function $x\mapsto 1/(1-\Lambda(x))$ is convex;
  3. (iii)
    the following statements are equivalent:
    1. (a)
      $T_{\Lambda}(a,x,X)$ is jointly convex in $(a,x)\in\overline{\mathbb{R}}\times\mathbb{R}$ for all $X\in L^{1}$;
    2. (b)
      $T_{\Lambda}(a,x,X)$ is jointly quasi-convex in $(a,x)\in\overline{\mathbb{R}}\times\mathbb{R}$ for all $X\in L^{1}$;
    3. (c)
      $T_{\Lambda}(a,x,X)$ is jointly convex in $(a,x,X)\in\overline{\mathbb{R}}\times\mathbb{R}\times L^{1}$;
    4. (d)
      $T_{\Lambda}(a,x,X)$ is jointly quasi-convex in $(a,x,X)\in\overline{\mathbb{R}}\times\mathbb{R}\times L^{1}$;
    5. (e)
      $\Lambda$ is constant on $\mathbb{R}$.
7 Conclusion
This paper introduces the $\Lambda$-Expected Shortfall ($\Lambda$-ES), a novel and theoretically grounded generalization of Expected Shortfall (ES) that robustly extends the $\Lambda$-VaR framework. We obtain an explicit representation of $\Lambda$-ES and verify that it satisfies several crucial properties as a desired counterpart to $\Lambda$-VaR. In particular, we show that $\Lambda$-ES is the smallest quasi-convex and law-invariant risk measure that dominates $\Lambda$-VaR. The dual representation of $\Lambda$-ES further connects it to established results on quasi-convex cash-subadditive risk measures. Our RU representation of $\Lambda$-ES provides useful insights for its potential applications to optimization problems. Practically, $\Lambda$-ES shares the advantages of $\Lambda$-VaR in the sense of flexible risk preferences and confidence levels, but more importantly, provides a useful alternative to $\Lambda$-VaR with additional benefits in risk management such as quasi-convexity and robustness.
The introduction of $\Lambda$-ES opens several promising avenues for future research. On the theoretical side, further investigation into its elicitability and connections to other axiomatically defined risk measures would lead to a deeper understanding of this new class of risk measures. Empirical studies applying $\Lambda$-ES in portfolio optimization, capital allocation, and stress testing with real-world financial data can be conducted for assessing its practical impact. Moreover, the development of efficient numerical algorithms for computing $\Lambda$-ES and solving $\Lambda$-ES-constrained optimization problems presents an important challenge for computational finance. $\Lambda$-ES holds substantial potential to enhance risk quantification and management practices, particularly in regulatory contexts that demand flexible yet robust risk measures.
Acknowledgments
RW acknowledges financial support from the Natural Sciences and Engineering Research Council (NSERC) of Canada (Grant Nos. RGPIN-2024-03728, CRC-2022-00141).
Appendix A Other possible formulations of Lambda ES
There may be many ways of generalizing ES using a parameter $\Lambda$. Below we explain a few possible ways of defining $\Lambda$-ES that fail to satisfy basic requirements, and thus they are not suitable definitions.
A.1 An algebraic formulation
We first consider an algebraic way of defining ES. We can rewrite $\mathrm{ES}_{\alpha}$ as (Acerbi and Tasche, 2002)
$\mathrm{ES}_{\alpha}(X)=\frac{1}{1-\alpha}\mathbb{E}\left[X\mathds{1}_{\{X\geqslant\mathrm{VaR}_{\alpha}(X)\}}\right]+\mathrm{VaR}_{\alpha}(X)(1-\alpha-\mathbb{P}(X\geqslant\mathrm{VaR}_{\alpha}(X))).$
Denote by $L^{1}_{c}$ the set of all random variables in $L^{1}$ with continuous distributions. We have
$\mathrm{ES}_{\alpha}(X)=\mathbb{E}\left[X\mid X\geqslant\mathrm{VaR}_{\alpha}(X)\right],~~~X\in L^{1}_{c}.$ (A.1)
To define $\Lambda$-ES on continuous random variables using the idea of formulation (A.1), a choice is:
$\rho_{\Lambda}(X)=\mathbb{E}\left[X\mid X\geqslant\mathrm{VaR}_{\Lambda}(X)\right],~~~X\in L^{1}_{c}.$ (A.2)
However, this formulation is not monotone, as the following example shows. Let $\Omega=[0,1]$ with $\mathbb{P}$ the Lebesgue measure. Let $0<\epsilon<10$. For $\omega\in\Omega$ define
$X(\omega)=\epsilon(\omega-0.1)+\mathds{1}_{(0.1,0.9]}(\omega)+10\cdot\mathds{1}_{(0.9,1]}(\omega)\mbox{~and~}Y(\omega)=\epsilon(\omega-1)+10\cdot\mathds{1}_{(0.9,1]}(\omega),$
so that $X\geqslant Y$. Let $\Lambda$ be a strictly decreasing function with $\Lambda(1)=0.1$ and $\Lambda(0)=0.9$. We can compute
$\mathrm{VaR}_{\Lambda}(X)=1\mathrm{~so~that~}\rho_{\Lambda}(X)=2+0.45\epsilon.$
On the other hand,
$\mathrm{VaR}_{\Lambda}(Y)=0\mathrm{~so~that~}\rho_{\Lambda}(Y)=10-0.05\epsilon.$
Taking $\epsilon\downarrow 0$ yields that $\rho_{\Lambda}$ in (A.2) is not monotone and is therefore undesirable.
A.2 A formulation based on the Rockafellar–Uryasev relation
Another possible formulation of $\Lambda$-ES is based on the RU relation in (3). Namely, for a decreasing function $\Lambda:\mathbb{R}\to(0,1)$, we may define the following candidate risk measure
$\rho_{\Lambda}(X)=\inf_{x\in\mathbb{R}}\left\{x+\frac{1}{1-\Lambda(x)}\mathbb{E}[(X-x)_{+}]\right\},~~~X\in L^{1}.$ (A.3)
Here, we use infimum because the minimum may not exist in general. Clearly, $\rho_{\Lambda}$ is monotone in both $\Lambda$ and $X$, is law-invariant, and specializes to $\mathrm{ES}_{\alpha}$ when $\Lambda\equiv\alpha$ for $\alpha\in(0,1)$. Comparing (A.3) to Theorem 4, we are optimizing (22) over the subset $\{(a,x)\in\mathbb{R}^{2}:a=x\}$ so that the optimum is larger. We have $\mathrm{ES}_{\Lambda}\leqslant\rho_{\Lambda}$ so that $\rho_{\Lambda}$ also dominates $\mathrm{VaR}_{\Lambda}$. However, the following counterexample shows that $\rho_{\Lambda}$ defined in (A.3) is not quasi-convex in general, and is thus not an ideal candidate for $\Lambda$-ES.
Let $a_{0},b_{0},\alpha,\beta,\epsilon\in\mathbb{R}$ with $b_{0}<a_{0}$, $\epsilon>0$, and $3/4<\beta<\alpha<1$. Let
$\Lambda_{0}(a)=\alpha+(\beta-\alpha)\mathds{1}_{\{a>a_{0}\}},~~a\in\mathbb{R}.$
It is clear that
$\rho_{\Lambda_{0}}(X)=\inf_{x\leqslant a_{0}}\left\{x+\frac{1}{1-\alpha}\mathbb{E}[(X-x)_{+}]\right\}\wedge\inf_{x>a_{0}}\left\{x+\frac{1}{1-\beta}\mathbb{E}[(X-x)_{+}]\right\},~~~X\in L^{1}.$
Take $Y,Z\in L^{\infty}$ such that
$\mathbb{P}(Z=a_{0}-\epsilon)=1-\mathbb{P}(Z=b_{0})=3/4,~Y=a_{0}+3\epsilon.$
It follows that
$\rho_{\Lambda_{0}}(Z)=\mathrm{ES}_{\alpha}(Z)=a_{0}-\epsilon,~\rho_{\Lambda_{0}}(Y)=a_{0}+3\epsilon,$
and $\rho_{\Lambda_{0}}\left(\frac{Y+Z}{2}\right)=a_{0}+\frac{1}{1-\beta}\mathbb{E}\left[\left(\frac{Y+Z}{2}-a_{0}\right)_{+}\right]=a_{0}+\frac{3/4}{1-\beta}\epsilon>a_{0}+3\epsilon.$
This indicates that $\rho_{\Lambda_{0}}$ is not quasi-convex.
A.3 A formulation based on the score function of Lambda VaR
Let $\Lambda:\mathbb{R}\to(0,1)$ be a decreasing function. As discussed in Section 5, a natural possible formulation of $\Lambda$-ES is
$\rho_{\Lambda}(X)=\min_{a\in\overline{\mathbb{R}}}\mathbb{E}[cS_{\Lambda}(a,X)+f(X)],~~X\in L^{\infty},$
where $c>0$ is a constant, $f:\mathbb{R}\to\mathbb{R}$ is a real function, and
$S_{\Lambda}(a,y)=(a-y)_{+}-\int_{y}^{a}\Lambda(t)\,\mathrm{d}t=(y-a)_{+}-\int_{a}^{y}(1-\Lambda(t))\,\mathrm{d}t,~~a\in\overline{\mathbb{R}},~y\in\mathbb{R},$
is the scoring function for $\mathrm{VaR}_{\Lambda}$ with
$\mathrm{VaR}_{\Lambda}(X)\in\operatorname*{arg\,min}_{a\in\overline{\mathbb{R}}}\mathbb{E}[S_{\Lambda}(a,X)],~~X\in L^{\infty}.$
The following argument shows that $\rho_{\Lambda}$ cannot satisfy quasi-convexity, normalization, and $\rho_{\Lambda}\geqslant\mathrm{VaR}_{\Lambda}$ simultaneously and is thus not a good candidate for $\Lambda$-ES.
  1. (i)
    Suppose that $\rho_{\Lambda}$ is normalized. For all $a\in\mathbb{R}$, $\mathrm{VaR}_{\Lambda}(a)=a$, and thus a=ρ_Λ(a)=cS_Λ(a,a)+f(a)=f(a). Therefore, we have $f(a)=a$ for all $a\in\mathbb{R}$.
  2. (ii)
    Suppose that $\rho_{\Lambda}$ is normalized and $\rho_{\Lambda}\geqslant\mathrm{VaR}_{\Lambda}$. It implies that $\rho_{\alpha^{*}}\geqslant\mathrm{VaR}_{\alpha^{*}}$ for all $\alpha^{*}\in[\inf_{x\in\mathbb{R}}\Lambda(x),\sup_{x\in\mathbb{R}}\Lambda(x)]$. For all $X\in L^{\infty}$,
    $\mathbb{E}\left[c(X-\mathrm{VaR}_{\alpha^{*}}(X))_{+}-c(1-\alpha^{*})(X-\mathrm{VaR}_{\alpha^{*}}(X))+X\right]$
    $=\rho_{\alpha^{*}}(X)\geqslant\mathrm{ES}_{\alpha^{*}}(X)$ [Theorem 5]
    $=\mathbb{E}\left[\frac{1}{1-\alpha^{*}}(X-\mathrm{VaR}_{\alpha^{*}}(X))_{+}-(X-\mathrm{VaR}_{\alpha^{*}}(X))+X\right]$ [by (3)]
    Therefore, $c\geqslant 1/(1-\alpha^{*})$ for all $\alpha^{*}\in[\inf_{x\in\mathbb{R}}\Lambda(x),\sup_{x\in\mathbb{R}}\Lambda(x)]$, and thus $c\geqslant 1/(1-\sup_{x\in\mathbb{R}}\Lambda(x))$.
  3. (iii)
[…]